### v1

# Robust Weight Initialization for Tanh Neural Networks with Fixed Point Analysis

Abstract
As a neural network’s depth increases, it can achieve strong generalization performance. Training, however, becomes challenging due to gradient issues. Theoretical research and various methods have been introduced to address this issues. However, research on weight initialization methods that can be effectively applied to tanh neural networks of varying sizes still needs to be completed. This paper presents a novel weight initialization method for Feedforward Neural Networks with tanh activation function. Based on an analysis of the fixed points of the function $\tanh(ax)$, our proposed method aims to determine values of $a$ that prevent the saturation of activations. A series of experiments on various classification datasets demonstrate that the proposed method is more robust to network size variations than the existing method. Furthermore, when applied to Physics-Informed Neural Networks, the method exhibits faster convergence and robustness to variations of the network size compared to Xavier initialization in problems of Partial Differential Equations.
1 Introduction
Deep learning has enabled substantial advancements in state-of-the-art performance across various domains (LeCun et al., 2015; He et al., 2016). In general, the expressivity of neural networks exponentially increases with depth (Poole et al., 2016; Raghu et al., 2017), enabling strong generalization performance. This increased depth, though, can result in vanishing or exploding gradients and poor signal propagation throughout the model (Bengio et al., 1993), prompting the development of various weight initialization methods. Xavier initialization (Glorot & Bengio, 2010) ensures signals stay in the non-saturated region for sigmoid and hyperbolic tangent activations, while He initialization (He et al., 2015) maintains stable variance for ReLU networks. Especially in ReLU neural networks, several weight initialization methods have been proposed to mitigate the dying ReLU problem, which hinders signal propagation in deep networks (Lu et al., 2019; Lee et al., 2024). However, to the best of our knowledge, research on the initialization method to tackle the stability of extremely deep tanh networks during training is still limited. Such networks commonly use Xavier initialization (Raissi et al., 2019; Jagtap et al., 2022; Rathore et al., 2024) and are widely applied in various domains, such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and Recurrent Neural Networks (RNNs) (Rumelhart et al., 1986), with performance often dependent on model size and initialization randomness (Liu et al., 2022).
The main contribution of this paper is the proposal of a simple weight initialization method for FeedForward Neural Networks (FFNNs) with tanh activation function. This method facilitates effective learning across a range of network sizes, outperforming Xavier initialization by reducing the need for extensive hyperparameter tuning such as the number of hidden layers and units. The theoretical foundation for this approach is provided through the fixed point of the function $\tanh(ax)$. We experimentally demonstrate that the proposed method achieves higher validation accuracy and lower validation loss compared to the Xavier initialization method across various FFNN network sizes on the MNIST, Fashion MNIST, and CIFAR-10 datasets. Additionally, the proposed method demonstrates its effectiveness in training across various network configurations within PINNs. Notably, while Xavier initialization shows decreasing loss as network depth increases, it fails to maintain performance beyond a certain depth, leading to increased loss and poor training outcomes. In contrast, the proposed method continues to improve performance even at greater depths, ensuring stable training and better results.
Contributions. Our contributions can be summarised as follows:
  • We show the conditions under which activation values do not vanish as we increase the depth of the neural network, using a fixed-point analysis (Section 3.1 and 3.2).
  • We propose a novel weight initialization method for tanh-based neural networks that has strong robustness to variations in network size (Section 3.2 and 3.3).
  • We experimentally demonstrate that the proposed method is more robust to variations in network size than Xavier initialization on image benchmark datasets and PINNs (Section 4).
2 Related works
The expressivity of neural networks typically grows exponentially with depth, resulting in improved generalization performance (Poole et al., 2016; Raghu et al., 2017). Weight initialization is crucial for training deep networks effectively (Saxe et al., 2014; Mishkin & Matas, 2016). Xavier (Glorot & Bengio, 2010) and He He et al. (2015) initialization are common initialization methods typically used with tanh and ReLU activation functions, respectively. Various initialization methods have been proposed to facilitate the training of deeper ReLU neural networks (Lu et al., 2019; Bachlechner et al., 2021; Zhao et al., 2022; Lee et al., 2024). However, to the best of our knowledge, research on weight initialization for neural networks with tanh activation remains limited. Tanh neural networks have been increasingly used, particularly in physics-informed neural networks (PINNs).
PINNs have shown promising results in solving forward, inverse, and multiphysics problems arising in science and engineering. (Lu et al., 2021; Karniadakis et al., 2021; Cuomo et al., 2022b; a; Yin et al., 2021; Wu et al., 2023; Hanna et al., 2022; Bararnia & Esmaeilpour, 2022; Shukla et al., 2020; Zhu et al., 2024; Hosseini et al., 2023; Mao et al., 2020). PINNs approximate solutions to partial differential equations (PDEs) using neural networks and are typically trained by minimizing a loss defined by the sum of least-squares that incorporates the residual of PDE, boundary conditions, and initial conditions. This loss is usually minimized using gradient-based optimizers such as Adam (Kingma, 2014), L-BFGS (Liu & Nocedal, 1989), or a combination of both. Universal approximation theories (Cybenko, 1989; Hornik et al., 1989; Hornik, 1991; Park et al., 2020; Guliyev & Ismailov, 2018b; Shen et al., 2022; Guliyev & Ismailov, 2018a; Maiorov & Pinkus, 1999; Yarotsky, 2017; Gripenberg, 2003) guarantee the capability and performance of neural networks as an approximation of the analytic solution to PDE. However, PINNs still face challenges in accuracy, stability, computational complexity, and tuning optimal hyperparameters of loss terms. To alleviate these issues, many authors have introduced enhanced versions of PINNs: (1) the self-adaptive loss balanced PINNs (lbPINNs) that automatically adjust the hyperparameters of loss terms during the training process (Xiang et al., 2022), (2) the Bayesian PINNs (B-PINNs) that are specialized to deal with forward and inverse nonlinear problems with noisy data (Yang et al., 2021), (3) Rectified PINNs (RPINNs) that are trained with the gradient information from the numerical solution by the multigrid method and designed for solving stationary PDEs  (Peng et al., 2022), (4) Auxiliary Pinns (A-PINNs) that effectively handle integro-differential equations (Yuan et al., 2022), (5) conservative PINNs (cPINNs) and exetended PINNs (XPINNs) that adopt the domain decomposition technique (Jagtap et al., 2020; Jagtap & Karniadakis, 2020), (6) parrel PINNs that reduces the computational cost of cPINNs and XPINNs (Shukla et al., 2021), (7) gradient-enhanced PINNs (gPINNs) that use the gradient of the PDE loss term with respect to the network inputs  (Yu et al., 2022).
PINNs primarily employ Xavier initialization for training (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023; Song et al., 2024), but our experimental results indicate that this method limits the performance of larger network sizes. Although there have been recent results on initialization methods for PINNs, most of them have relied on transfer learning (Tarbiyati & Nemati Saray, 2023). Thus, we propose a weight initialization method that does not require transfer learning and is robust to variations in network size.
(a) Xavier Initialization
(b) Proposed Initialization
Figure 1: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using Xavier initialization (Left) and the proposed initialization (Right). Experiments were conducted on networks with 10,000 hidden layers, each having the same number of nodes: $16$, $32$, $64$, or $128$.
3 Proposed Weight Initialization method
In this section, we discuss the proposed weight initialization method. Subsection 3.1 introduces the theoretical motivation behind the methodology. Subsection 3.2 presents how to derive the initial weight matrix that satisfies the conditions outlined in Subsection 3.1. Finally, in Subsection 3.3, we suggest the optimal hyperparameter $\sigma_{z}$ in the proposed method.
3.1 Theoretical motivation
Experimental results in Figure 1 reveal that when Xavier initialization is employed in FFNNs with tanh activation, the distribution of activation values tends to cluster around zero in deeper layers. This vanishing of activation values can hinder the training process due to a discrepancy between the activation values and the desired output. However, theoretically preventing this phenomena is not straightforward. In this subsection, we gives a theoretical analysis based on a fixed point of $\tanh(ax)$ to bypass the phenomena. Before giving the theoretical foundations, consider the basic results for a tanh activation function. Recall that $x^{*}$ is a fixed point of a function $f$ if $x^{*}$ belongs to both the domain and the codomain of $f$, and $f(x^{*})=x^{*}$. The proofs of Lemma 1 and Lemma 2 are provided in Appendix A.
Lemma 1.
For a fixed $a>0$ define the function $\phi_{a}:\mathbb{R}\to\mathbb{R}$ given as
$\phi_{a}(x):=\tanh(ax).$
Then, there exists a fixed point $x^{\ast}$. Furthermore,
  • $(1)$
    if $0<a\leq 1$, then $\phi$ has a unique fixed point $x^{\ast}=0$.
  • $(2)$
    if $a>1$, then $\phi$ has three distinct fixed points: $x^{\ast}=-\xi_{a}$, 00, $\xi_{a}$ such that $\xi_{a}>0$.
Remark that the function $\phi_{a}$ can be considered as one-layer tanh FFNN.
Lemma 2.
For a given initial value $x_{0}>0$ define
$x_{n+1}=\phi_{a}(x_{n}),\quad n=0,1,2,\ldots.$
Then $\{x_{n}\}_{n=1}^{\infty}$ converges regardless of the positive initial value $x_{0}>0$. Moreover,
  • $(1)$
    if $0<a\leq 1$, then $x_{n}\to 0$ as $n\to\infty$.
  • $(2)$
    if $a>1$, then $x_{n}\to\xi_{a}$ as $n\to\infty$.
Note that the parameter $a$ in Lemma 2 does not change across all iterations. In Propositions 3 and Corollary 4, we address cases where the value of $a$ varies with each iteration.
Proposition 3.
Let $\{a_{n}\}_{n=1}^{\infty}$ be a positive real sequence, i.e., $a_{n}>0$ for all $n\in\mathbb{N}$, such that only finitely many elements are greater than $1$. Suppose that $\{\Phi_{m}\}_{m=1}^{\infty}$ is a sequence of functions defined as for each $m\in\mathbb{N}$
$\Phi_{m}=\phi_{a_{m}}\circ\phi_{a_{m-1}}\circ\cdots\circ\phi_{a_{1}}.$
Then for any $x\in\mathbb{R}$
$\lim_{m\to\infty}\Phi_{m}(x)=0.$
Proof.
Set $N=\max\{n|a_{n}>1\}$. Define the sequences $\{b_{n}\}_{n=1}^{\infty}$ and $\{c_{n}\}_{n=1}^{\infty}$ such that $b_{n}=c_{n}=a_{n}$ for $n\leq N$, with $b_{n}=0$ and $c_{n}=1$ for $n>N$. Suppose that $\{\hat{\Phi}_{m}\}_{m=1}^{\infty}$ and $\{\tilde{\Phi}_{m}\}_{m=1}^{\infty}$ are sequences of functions defined as for each $m\in\mathbb{N}$
$\hat{\Phi}_{m}=\phi_{b_{m}}\circ\phi_{b_{m-1}}\circ\cdots\circ\phi_{b_{1}},\quad\tilde{\Phi}_{m}=\phi_{c_{m}}\circ\phi_{c_{m-1}}\circ\cdots\circ\phi_{c_{1}}.$
Then, the inequality $\hat{\Phi}_{m}\leq\Phi_{m}\leq\tilde{\Phi}_{m}$ holds for all $m$. By Lemma 1, for any $x\geq 0$, we have $\lim_{m\to\infty}\hat{\Phi}_{m}=0$ and $\lim_{m\to\infty}\tilde{\Phi}_{m}=0$. Therefore, the Squeeze Theorem guarantees that $\lim_{m\to\infty}\Phi_{m}(x)=0$.
Corollary 4.
Let $\epsilon>0$ be given. Suppose that $\{a_{n}\}_{n=1}^{\infty}$ be a positive real sequence such that only finitely many elements are lower than $1+\epsilon$. Then for any $x\in\mathbb{R}\setminus\{0\}$
$\left|\lim_{m\to\infty}\Phi_{m}(x)\right|\geq\xi_{1+\epsilon}$
Proof.
Set $N=\max\{n\mid a_{n}<1+\epsilon\}$. Define the sequence $\{b_{n}\}_{n=1}^{\infty}$ such that $b_{n}=a_{n}$ for $n\leq N$, and $b_{n}=1+\epsilon$ for $n>N$. The remainder of the proof is analogous to the proof of Proposition 3. ∎
3.2 The derivation of the proposed weight initialization method
To establish the notation, consider a feedforward neural network with $L$ layers. The network processes $K$ training samples, denoted as pairs $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{K}$, where $\bm{x}_{i}\in\mathbb{R}^{N_{x}}$ is training input and $\bm{y}_{i}\in\mathbb{R}^{N_{y}}$ is its corresponding output. The iterative computation at each layer $\ell$ is defined as follows:
$\bm{x}^{\ell}=\tanh(\mathbf{W}^{\ell}\bm{x}^{{\ell}-1}+\mathbf{b}^{\ell})\in\mathbb{R}^{N_{\ell}}\quad\text{for all }\ell=1,\ldots,L,$
where $\mathbf{W}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ is the weight matrix, $\mathbf{b}^{\ell}\in\mathbb{R}^{N_{\ell}}$ is the bias, and $\tanh(\cdot)$ is an element-wise activation hyperbolic tangent function.
We present a simplified analysis of signal propagation in FFNNs with the tanh activation function. For notational convenience, it is assumed that all hidden layers, as well as the input and output layers, have a dimension of $n$, i.e., $N_{\ell}=n$ for all $\ell$. Given an arbitrary input vector $\bm{x}=(x_{1},\ldots,x_{n})$, the first layer activation $\bm{x}^{1}=\tanh(\mathbf{W}^{1}\bm{x})$ can be expressed component-wise as:
$x^{1}_{i}=\tanh\big{(}w^{1}_{i1}x_{1}+\cdots+w^{1}_{in}x_{n}\big{)}=\tanh\bigg{(}\Big{(}w^{1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{1}_{ij}x_{j}}{x_{i}}\Big{)}x_{i}\bigg{)},\,\text{for }i=1,\dots,n.$
For the $k+1$-th layer, $i=1,\ldots,n$, this expression can be generalized as:
$x_{i}^{k+1}=\tanh\left(a_{i}^{k+1}x_{i}^{k}\right),\,\text{where}\quad a_{i}^{k+1}=w^{k+1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{k+1}_{ij}x_{j}^{k}}{x_{i}^{k}}.$ (1)
According to Lemma 2, when $a>1$, for an arbitrary initial value $x_{0}>0$ or $x_{0}<0$, the sequence $\{x_{k}\}$ defined by $x_{k+1}=\tanh(ax_{k})$ converges to $\xi_{a}$ or $-\xi_{a}$, respectively, as $k\to\infty$. This result indicates that the sequence converges to the fixed point $\xi_{a}$ regardless of the initial value $x_{0}$ and ensures that the activation values do not vanish as network depth increases. Furthermore, by Lemma 2, if $a^{k}_{i}\leq 1$ for all $N\leq k\leq L$, then $x_{i}^{L}$ approaches zero. Therefore, to ensure that ((((i)))) $a_{i}^{k}$ remains close to 1 and ((((ii)))) $a_{i}^{k}\leq 1$ does not hold for all $N\leq k\leq L$, we design the initial weight matrix as $\mathbf{W}^{\ell}=\mathbf{D}^{\ell}+\mathbf{Z}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$, where $\mathbf{D}^{\ell}_{i,j}=1$ if $i\equiv j\pmod{N_{\ell-1}}$, 0 otherwise, and $\mathbf{Z}^{\ell}$ is a noise matrix drawn from $\mathcal{N}(0,\sigma_{z}^{2})$, where $\sigma_{z}$ is set to $\alpha/\sqrt{N^{\ell-1}}$ and $\alpha=0.085$. Then $a_{i}^{k+1}$ follows the distribution:
$a_{i}^{k+1}\sim\mathcal{N}\Bigg{(}1,\,\sigma_{z}^{2}+\sigma_{z}^{2}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\bigg{(}\frac{x_{j}^{k}}{x_{i}^{k}}\bigg{)}^{2}\Bigg{)}.$ (2)
According to Equation 2, the mean of $a_{i}^{k+1}$ is $1$, so choosing an appropriate $\sigma_{z}$ satisfies condition ((((i)))). For condition ((((ii)))), if $x_{i}^{k}$ becomes small relative to other elements in $\bm{x}^{k}$, the variance of $a_{i}^{k+1}$ increases, as indicated by Equation 2. As a result, the probability that the absolute value of $x_{i}^{k+1}$ surpasses that of $x_{i}^{k}$ is higher. However, if $\sigma_{z}$ is too small, the increase in the variance of $a_{i}^{k+1}$ becomes limited. Therefore, choosing an appropriate $\sigma_{z}$ is crucial.
(a) Same dimension
(b) Varying dimensions
Figure 2: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using the proposed initialization with $\alpha$ set to $0.04,0.085,0.15,$ and $0.5$. Network with $10,000$ hidden layers, each with 32 nodes (Left), and a network with alternating hidden layers of $64$ and $32$ nodes (Right).
3.3 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
In this subsection, we discuss how $\sigma_{z}$ impacts the scale of the activation values. Equation 2 indicates that $a_{i}^{k}$ follows a normal distribution, with variance depending on $\sigma_{z}$. Firstly, we experimentally investigated the impact of $\sigma_{z}$ on the scale of the activation values. As demonstrated in Figure 2, increasing $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ causes the activation values in any layer to be distributed over a broader range. However, setting $\sigma_{z}$ to a large value can lead to saturation, where most activations converge towards $-1$ and $1$. If $\sigma_{z}$ is too large, the probability that $a_{i}^{(k)}$ takes values far from $1$ (e.g., -10, 5, etc.) increases. This, in turn, increases the value of $1+\epsilon$ mentioned in Corollary 4, potentially bounding the activation values in sufficiently deep layers by $\xi_{1+\epsilon}$. Consequently, the activation values in deeper layers become less likely to approach zero and tend to saturate toward specific values. On the other hand, if $\sigma_{z}$ is too small, as mentioned in Subsection 3.2, the variance of $a_{i}^{k}$ becomes restricted. This is demonstrated experimentally in Figure 2, when $\alpha=0.00001$. For this reason, we experimentally found an optimal $\sigma_{z}=\alpha/\sqrt{N^{\ell-1}}$, with $\alpha=0.085$, that is neither too large nor too small. Results from experiments solving the Burgers’ equation using PINNs with varying $\sigma_{z}$ are presented in Appendix B.1.
Table 1: Validation accuracy and loss are presented for FFNNs with varying numbers of nodes $(2,8,32,128)$, each with $20$ hidden layers using the tanh activation function. All models were trained for $20$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented. When comparing different initialization methods under the same experimental settings, the better-performing method is highlighted in bold. Underlined values indicate the highest accuracy when only the number of nodes is varied.
2 8 32 128
MNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 49.78 1.632 68 0.958 91.67 0.277 95.45 0.154
Proposed 62.82 1.185 77.95 0.706 92.51 0.255 96.12 0.134
FMNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 42.89 1.559 68.55 0.890 81.03 0.533 86.2 0.389
Proposed 51.65 1.324 71.31 0.777 83.06 0.475 87.12 0.359
CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 32.82 1.921 43.51 1.608 48.62 1.473 47.58 1.510
Proposed 38.16 1.780 47.04 1.505 48.80 1.463 48.51 1.471
Table 2: Validation accuracy and loss are presented for FFNNs with varying numbers of layers $(10,50,100)$, each with $64$ number of nodes using the tanh activation function. All models were trained for $40$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented.
10 50 100
MNIST Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 96.55 0.112 96.57 0.123 94.08 0.194
Proposed 97.04 0.102 96.72 0.109 96.06 0.132
FMNIST Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 88.73 0.319 87.72 0.344 83.41 0.463
Proposed 89.42 0.305 88.51 0.324 86.01 0.382
CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss
Xavier 48.39 1.468 47.87 1.474 46.71 1.503
Proposed 48.41 1.458 48.71 1.461 48.96 1.437
4 Experiments
In this section, we conduct a series of experiments to validate the proposed weight initialization method. In Subsection 4.1, we evaluate the performance of an FFNN with the tanh activation function on benchmark datasets. In Subsection 4.2, we solve the Burgers’ equation and Allen-Cahn equation using Physics-Informed Neural Networks. Both experiments are conducted across various network sizes to verify whether the proposed method consistently performs well, independent of network depth and width. The experiments were conducted in TensorFlow without skip connections, normalization layers, and learning rate decay in any of the experiments.
4.1 Width Independence in Classification Task
To evaluate the effectiveness of the proposed weight initialization method, we conduct experiments on the MNIST, Fashion MNIST, and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, utilizing the Adam optimizer. All experiments are conducted with a batch size of 64 and a learning rate of 0.0001. Fifteen percent of the total dataset is allocated for validation.
We apply the proposed weight initialization method to evaluate its effectiveness in training tanh FFNNs, emphasizing its robustness to variations in network width. Four tanh FFNNs are created, each with 20 hidden layers, and with 2, 8, 32, and 128 nodes per hidden layer, respectively. In Table 1, for both the MNIST and Fashion MNIST datasets, the network with 128 nodes achieves the highest accuracy and lowest loss when our proposed method is employed. However, for the CIFAR-10 dataset, the network with 32 nodes yields the highest accuracy and lowest loss when employing the proposed method. In summary, our proposed method demonstrates robustness regardless of the number of nodes in tanh FFNNs. We provide more detailed experimental results in Appendix B.2.
4.2 Depth Independence in Classification Task
It is well known that the expressivity of neural networks generally increases exponentially with depth, enabling strong generalization performance (Poole et al., 2016; Raghu et al., 2017). Therefore, we employ the proposed weight initialization method to investigate its effectiveness in training deep FFNNs with the tanh activation function, emphasizing its robustness to variations in network depth. We create three tanh FFNNs, each with 64 nodes in all hidden layers, but with 10, 50, and 100 hidden layers, respectively. In Table 2, for both the MNIST and Fashion MNIST datasets, the network with 10 hidden layers achieves the highest accuracy and lowest loss when our proposed method is employed. Both initialization methods perform best in networks with the fewest layers, with performance degrading as the depth increases. However, for the CIFAR-10 dataset, we observe that the performance of the proposed method improves as the number of layers increases.
Furthermore, we conduct experiments with varying hidden layer dimensions, as shown in Figure 3. The network consists of 60 hidden layers, where the number of nodes alternates between 32 and 16 in each layer. We demonstrate superior performance in terms of both loss and accuracy across all epochs on the MNIST and CIFAR-10 datasets.
(a) MNIST Accuracy
(b) MNIST Loss
(c) CIFAR10 Accuracy
(d) CIFAR10 Loss
Figure 3: Validation accuracy and loss for a tanh FFNN with 60 hidden layers, where the number of nodes alternates between 32 and 16 across layers, repeated 30 times. The model was trained for 20 epochs on the MNIST and CIFAR-10 datasets.
4.3 Network size independence in pinn
Xavier initialization is the primary method used for training PINNs  (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023). In this section, we experimentally demonstrate that the method’s training performance is highly dependent on randomness and network size. Additionally, empirical results are provided demonstrating that the proposed method is more robust to variations in network size.
All experiments on Physics-Informed Neural Networks (PINNs) use full-batch training with a learning rate of 0.001. In this section, we solve the Allen-Cahn and Burgers’ equations using a tanh FFNN-based PINN with 20,000 collocation points. For the Allen-Cahn equation, the diffusion coefficient is set to $d=0.01$. The initial condition is defined as $u(x,0)=x^{2}\cos(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=-1$ and $u(1,t)=-1$, applied over the time interval $t\in[0,1]$. Similarly, for the Burgers’ equation, a viscosity coefficient of $\nu=0.01$ is employed. The initial condition is given by $u(x,0)=-\sin(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=0$ and $u(1,t)=0$ imposed for $t\in[0,1]$.
The Allen-Cahn equation is expressed as:
$\frac{\partial u}{\partial t}-d\frac{\partial^{2}u}{\partial x^{2}}=-\frac{u^{3}+u}{d}$
where $u(x,t)$ represents the solution, $d$ is the diffusion coefficient, and the nonlinear term $u^{3}-u$ models the phase separation dynamics.
The Burgers’ equation is given by:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}}$
where $u(x,t)$ is the velocity field, and $\nu$ is the viscosity coefficient.
Eight tanh FFNNs are created, each with 16 nodes in all hidden layers, but with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively. As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 20. However, as the depth increases, the loss gradually rises. In contrast, the proposed method achieves the lowest loss at a depth of 50 and maintains a loss of 0.00057 even at a depth of 80 layers. For the Burgers’ equation, the proposed method achieves the lowest loss at a depth of 60, while at the same depth, Xavier initialization results in a loss that is an order of magnitude higher (approximately $10^{2}$ difference).
Next, we double the number of nodes to observe the impact of node size on the loss. Eight new tanh FFNNs are created, each with 32 nodes in all hidden layers, and with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively. As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 30. However, as the depth increases, the model becomes untrainable, with a loss of 0.694 at a depth of 80. In contrast, the proposed method achieves the lowest loss at a depth of 40 and maintains a loss of 0.00059 even at a depth of 80 layers. For the Burgers’ equation, both methods show similar loss values up to a depth of 30. Beyond a depth of 40, however, the loss steadily increases with the Xavier method, while the proposed method records the lowest loss at a depth of 50.
Table 3: A PINN loss is presented for FFNNs with varying numbers of layers $(5,10,20,30,40,50,60,80)$ using the tanh activation function. The top table shows results with $16$ nodes per layer, and the bottom table shows results with $32$ nodes per layer. All models were trained for $300$ iterations using Adam and $300$ iterations using L-BFGS. The median PINN loss from the final iteration for the Burgers and Allen–Cahn equations, computed over 5 runs, is presented.
Allen-Cahn (16 Nodes) 5 10 20 30 40 50 60 80
Xavier 9.58e-04 8.16e-04 7.61e-04 1.06e-03 1.1e-03 1.24e-03 3.55e-03 1.81e-03
Proposed 9.21e-04 7.29e-04 5.76e-04 5.29e-04 5.37e-04 4.03e-04 4.73e-04 5.77e-04
Burgers (16 Nodes) 5 10 20 30 40 50 60 80
Xavier 6.97e-03 1.11e-02 7.9e-03 9.71e-03 2.45e-02 2.65e-02 6.5e-02 5.71e-02
Proposed 6.19e-03 5.08e-03 5.28e-03 9.31e-04 3.56e-03 8.27e-04 3.43e-04 2.05e-03
Allen-Cahn (32 Nodes) 5 10 20 30 40 50 60 80
Xavier 3.13e-01 5.03e-02 3.64e-03 2.37e-03 4.03e-03 5.27e-03 1.73e-02 6.94e-01
Proposed 1.04e-03 6.92e-04 5.34e-04 4.26e-04 3.31e-04 3.52e-04 3.85e-04 5.96e-04
Burgers (32 Nodes) 5 10 20 30 40 50 60 80
Xavier 1.12e-02 3.53e-03 2.72e-03 1.81e-03 7.60e-03 8.56e-03 9.86e-03 1.66e-01
Proposed 4.14e-03 4.11e-03 1.58e-03 1.29e-03 7.96e-04 5.85e-04 9.80e-04 1.47e-03
5 Conclusion
In this paper, we have introduced a novel weight initialization method for tanh FFNNs, grounded in the theoretical analysis of fixed points of the $\tanh(ax)$ function. Through our fixed-point analysis, we established conditions under which the vanishing or exploding of activation values can be prevented, even as the depth of the network increases.
Our proposed method exhibits strong robustness to variations in network size, as demonstrated across a variety of FFNN configurations and benchmark datasets, including MNIST, Fashion MNIST, and CIFAR-10. In contrast to Xavier initialization, which struggles to maintain stable performance as network depth increases, the proposed method consistently achieves superior results by preserving activation values. Furthermore, we explored the impact of the initialization hyperparameter $\sigma_{z}$ on the distribution of activation values. We demonstrated both theoretically and experimentally that the choice of $\sigma_{z}$ plays a significant role in maintaining the proper range of activations, balancing between vanishing and saturation. In the context of PINNs, the proposed initialization method shows improved performance in solving PDEs such as the Burgers’ equation and the Allen-Cahn equation. By maintaining a stable loss function and achieving faster convergence compared to Xavier initialization, our method demonstrates its practical utility in training networks for physical systems.
A key advantage of the proposed method lies in its robustness to network depth and width, significantly reducing the need for extensive hyperparameter tuning. By maintaining stable performance across varying network configurations, our approach helps to minimize the time and effort spent on searching for optimal network architectures, allowing researchers to focus on model design and other aspects of the training process. This makes the proposed method particularly valuable in large-scale and resource-constrained applications where efficient training is critical.
Appendix A Proofs of the theoretical results
A.1 Proof of Lemma 1
Proof.
We define $g(x)=\tanh(ax)-x$. Since $g(x)$ is continuous, and $g(-M)>0$, $g(M)<0$ for a large real number $M\in\mathbb{R}^{+}$, the Intermediate Value Theorem guarantees the existence of a point $x$ such that $g(x)=0$.
First, consider the case $0<a\leq 1$. Since $0<a\leq 1$, the derivative $g^{\prime}(x)=a\cdot\text{sech}^{2}(ax)-1$ satisfies $-1\leq g^{\prime}(x)\leq a-1<0$ for all $x$. Hence, $g(x)$ is strictly decreasing and therefore $g(x)$ has the unique root. At $x=0$, $\phi(0)=\tanh(a\cdot 0)=0$. Hence, $x=0$ is the unique fixed point.
Let us consider the case $a>1$. For $0<x\ll 1$, $\tanh(ax)-x\approx(a-1)x$. Since $a>1,\tanh(ax)-x>0$. On the other hand, since $|\tanh(ax)|<1$ for all $x$,
$\lim_{x\to\infty}[-1-x]\leq\lim_{x\to\infty}[\tanh(ax)-x]\leq\lim_{x\to\infty}[1-x].$
By the squeeze theorem, $\lim_{x\to\infty}[\tanh(ax)-x]=-\infty$. By the intermediate value theorem, therefore, there exists at least one $x>0$ such that $\tanh(ax)=x$. To establish the uniqueness of the positive fixed point, we investigate the derivative $g^{\prime}(x)=a\operatorname{sech}^{2}(ax)-1$. We find the critical points to be $x=\pm\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}})$. It is straightforward to see that $g^{\prime}(x)>0$ in $\left(-\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}}),\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}})\right)$ and $g^{\prime}(x)<0$ in \$\mathbb{R}\backslash\left(-\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}}),\frac{1}{a}\sec^{-1}(\frac{1}{\sqrt{a}})\right)$. i.e. $g(x)=0$ has exactly two fixed points. Because $g(x)$ is an odd function, if $x^{\ast}$ is a solution, then $-x^{\ast}$ is also a solution. Thus, for $a>1$, there exists a unique positive fixed point if $x>0$ and a unique negative fixed point if $x<0$. ∎
A.2 Proof of Lemma 2
Proof.
(1) Since $(\tanh(ax))^{\prime}=a\operatorname{sech}^{2}(ax)<1$ for all $x>0$, it holds that $x_{n+1}=\phi_{a}(x_{n})<x_{n}$ for all $n\in\mathbb{N}$. Thus the sequence $\{x_{n}\}_{n=1}^{\infty}$ is decreasing. Since $x_{n}>0$ for all $n\in\mathbb{N}$, by the monotone convergence theorem, it converges to the fixed point $x^{*}=0$.
(2) Let $x_{0}<\xi_{a}$. Since $\phi^{\prime}(x)$ decreasing for $x\geq 0$, with $\phi^{\prime}(0)>1$ and $\xi_{a}$ is the unique fixed point for $x>0$, it holds that $x_{n}<x_{n+1}<\xi_{a}$ for all $n\in\mathbb{N}$. Thus, by the monotone convergence theorem, the sequence converges to the fixed point $\xi_{a}$. The proof is similar when $x_{0}>\xi_{a}$. ∎
Appendix B Additional experimental results
B.1 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
(a) Same dimension
(b) Varying dimensions
Figure 4: Here, ’STD’ refers to $\sigma_{z}$. (a) shows the PINN loss for the Burgers’ equation, using an FFNN with 30 layers and 32 nodes in each hidden layer. (b) shows the PINN loss for an FFNN with 30 layers, where the hidden layers alternate between 64 and 32 nodes, repeated 15 times. Each experiment was repeated 10 times with different random seeds.
B.2 Width independence in Classification tasks
Figure 5: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the MNIST dataset, with 10 different random seeds.
Figure 6: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the Fashion MNIST dataset, with 10 different random seeds.
Figure 7: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers. All models were trained for 20 epochs on the CIFAR-10 dataset, with 10 different random seeds.

### v2

@@
 Abstract
-As a neural network’s depth increases, it can achieve strong generalization performance.
-Training, however, becomes challenging due to gradient issues.
-Theoretical research and various methods have been introduced to address this issues.
-However, research on weight initialization methods that can be effectively applied to tanh neural networks of varying sizes still needs to be completed.
-This paper presents a novel weight initialization method for Feedforward Neural Networks with tanh activation function.
-Based on an analysis of the fixed points of the function $\tanh(ax)$, our proposed method aims to determine values of $a$ that prevent the saturation of activations.
-A series of experiments on various classification datasets demonstrate that the proposed method is more robust to network size variations than the existing method.
-Furthermore, when applied to Physics-Informed Neural Networks, the method exhibits faster convergence and robustness to variations of the network size compared to Xavier initialization in problems of Partial Differential Equations.
+As a neural network’s depth increases, it can improve generalization performance.
+However, training deep networks is challenging due to gradient and signal propagation issues.
+To address these challenges, extensive theoretical research and various methods have been introduced.
+Despite these advances, effective weight initialization methods for tanh neural networks remain insufficiently investigated.
+This paper presents a novel weight initialization method for neural networks with tanh activation function.
+Based on an analysis of the fixed points of the function $\tanh(ax)$, the proposed method aims to determine values of $a$ that mitigate activation saturation.
+A series of experiments on various classification datasets and physics-informed neural networks demonstrates that the proposed method outperforms Xavier initialization methods (with or without normalization) in terms of robustness across different network sizes, data efficiency, and convergence speed.
+Code is available at https://github.com/1HyunwooLee/Tanh-Init.
 1 Introduction
-Deep learning has enabled substantial advancements in state-of-the-art performance across various domains (LeCun et al., 2015; He et al., 2016).
-In general, the expressivity of neural networks exponentially increases with depth (Poole et al., 2016; Raghu et al., 2017), enabling strong generalization performance.
-This increased depth, though, can result in vanishing or exploding gradients and poor signal propagation throughout the model (Bengio et al., 1993), prompting the development of various weight initialization methods.
-Xavier initialization (Glorot & Bengio, 2010) ensures signals stay in the non-saturated region for sigmoid and hyperbolic tangent activations, while He initialization (He et al., 2015) maintains stable variance for ReLU networks.
+Deep learning has significantly advanced state-of-the-art performance across various domains (LeCun et al., 2015; He et al., 2016).
+The expressivity of neural networks increases exponentially with depth, enhancing generalization performance (Poole et al., 2016; Raghu et al., 2017).
+However, deeper networks often face gradient and signal propagation issues (Bengio et al., 1993).
+These challenges have driven the development of effective weight initialization methods designed for various activation functions.
+Xavier initialization (Glorot & Bengio, 2010) prevents saturation in sigmoid and tanh activations, while He initialization (He et al., 2015) stabilizes variance for ReLU networks.
 Especially in ReLU neural networks, several weight initialization methods have been proposed to mitigate the dying ReLU problem, which hinders signal propagation in deep networks (Lu et al., 2019; Lee et al., 2024).
-However, to the best of our knowledge, research on the initialization method to tackle the stability of extremely deep tanh networks during training is still limited.
-Such networks commonly use Xavier initialization (Raissi et al., 2019; Jagtap et al., 2022; Rathore et al., 2024) and are widely applied in various domains, such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and Recurrent Neural Networks (RNNs) (Rumelhart et al., 1986), with performance often dependent on model size and initialization randomness (Liu et al., 2022).
+However, to the best of our knowledge, research on initialization methods that are robust across different sizes of tanh networks is underexplored.
+Tanh networks commonly employ Xavier initialization (Raissi et al., 2019; Jagtap et al., 2022; Rathore et al., 2024) and are applied in various domains, such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and Recurrent Neural Networks (RNNs) (Rumelhart et al., 1986), with performance often dependent on model size and initialization randomness (Liu et al., 2022).
 The main contribution of this paper is the proposal of a simple weight initialization method for FeedForward Neural Networks (FFNNs) with tanh activation function.
-This method facilitates effective learning across a range of network sizes, outperforming Xavier initialization by reducing the need for extensive hyperparameter tuning such as the number of hidden layers and units.
-The theoretical foundation for this approach is provided through the fixed point of the function $\tanh(ax)$.
-We experimentally demonstrate that the proposed method achieves higher validation accuracy and lower validation loss compared to the Xavier initialization method across various FFNN network sizes on the MNIST, Fashion MNIST, and CIFAR-10 datasets.
-Additionally, the proposed method demonstrates its effectiveness in training across various network configurations within PINNs.
-Notably, while Xavier initialization shows decreasing loss as network depth increases, it fails to maintain performance beyond a certain depth, leading to increased loss and poor training outcomes.
-In contrast, the proposed method continues to improve performance even at greater depths, ensuring stable training and better results.
-Contributions. Our contributions can be summarised as follows:
-  • We show the conditions under which activation values do not vanish as we increase the depth of the neural network, using a fixed-point analysis (Section 3.1 and 3.2).
-  • We propose a novel weight initialization method for tanh-based neural networks that has strong robustness to variations in network size (Section 3.2 and 3.3).
-  • We experimentally demonstrate that the proposed method is more robust to variations in network size than Xavier initialization on image benchmark datasets and PINNs (Section 4).
+The proposed method is data-efficient and demonstrates robustness across different network sizes.
+Moreover, it reduces dependency on normalization techniques such as Batch Normalization (BN) (Ioffe, 2015) and Layer Normalization (LN) (Ba, 2016).
+As a result, it alleviates the need for extensive hyperparameter tuning, such as selecting the number of hidden layers and units, while also eliminating the computational overhead associated with normalization.
+The theoretical foundation of this approach is based on the fixed point of the function $\tanh(ax)$.
+We evaluate the proposed method on two tasks: classification and PINNs.
+For classification tasks, we assess its performance across various FFNN sizes using standard benchmark datasets.
+The results demonstrate improved validation accuracy and lower loss compared to Xavier initialization with BN or LN.
+For PINNs, the method exhibits robustness across diverse network sizes and demonstrates its effectiveness in solving a wide range of PDE problems.
+Notably, for both tasks, the proposed method outperforms Xavier initialization in data efficiency; that is, it achieves improved performance even with limited data.
+Our main contributions can be summarized as follows:
+  • We identify the conditions under which activation values do not vanish as the neural network depth increases, using a fixed-point analysis (Section 3.1 and 3.2).
+  • We propose a novel weight initialization method for tanh neural networks that is robust across different network sizes and demonstrates high data efficiency (Section 3.2 and 3.3).
+  • We experimentally show that the proposed method is more robust across different network sizes on image benchmarks and PINNs (Section 4).
+  • We experimentally show that the proposed method is more data-efficient than Xavier initialization, with or without normalization, on image benchmarks and PINNs (Section 4).
 2 Related works
-The expressivity of neural networks typically grows exponentially with depth, resulting in improved generalization performance (Poole et al., 2016; Raghu et al., 2017).
-Weight initialization is crucial for training deep networks effectively (Saxe et al., 2014; Mishkin & Matas, 2016).
+The expressivity of neural networks grows exponentially with depth, resulting in improved generalization performance (Poole et al., 2016; Raghu et al., 2017).
+Proper weight initialization is crucial for effectively training deep networks (Saxe et al., 2014; Mishkin & Matas, 2016).
 Xavier (Glorot & Bengio, 2010) and He He et al.
 (2015) initialization are common initialization methods typically used with tanh and ReLU activation functions, respectively.
-Various initialization methods have been proposed to facilitate the training of deeper ReLU neural networks (Lu et al., 2019; Bachlechner et al., 2021; Zhao et al., 2022; Lee et al., 2024).
-However, to the best of our knowledge, research on weight initialization for neural networks with tanh activation remains limited.
-Tanh neural networks have been increasingly used, particularly in physics-informed neural networks (PINNs).
-PINNs have shown promising results in solving forward, inverse, and multiphysics problems arising in science and engineering.
-(Lu et al., 2021; Karniadakis et al., 2021; Cuomo et al., 2022b; a; Yin et al., 2021; Wu et al., 2023; Hanna et al., 2022; Bararnia & Esmaeilpour, 2022; Shukla et al., 2020; Zhu et al., 2024; Hosseini et al., 2023; Mao et al., 2020).
-PINNs approximate solutions to partial differential equations (PDEs) using neural networks and are typically trained by minimizing a loss defined by the sum of least-squares that incorporates the residual of PDE, boundary conditions, and initial conditions.
-This loss is usually minimized using gradient-based optimizers such as Adam (Kingma, 2014), L-BFGS (Liu & Nocedal, 1989), or a combination of both.
-Universal approximation theories (Cybenko, 1989; Hornik et al., 1989; Hornik, 1991; Park et al., 2020; Guliyev & Ismailov, 2018b; Shen et al., 2022; Guliyev & Ismailov, 2018a; Maiorov & Pinkus, 1999; Yarotsky, 2017; Gripenberg, 2003) guarantee the capability and performance of neural networks as an approximation of the analytic solution to PDE.
-However, PINNs still face challenges in accuracy, stability, computational complexity, and tuning optimal hyperparameters of loss terms.
-To alleviate these issues, many authors have introduced enhanced versions of PINNs: (1) the self-adaptive loss balanced PINNs (lbPINNs) that automatically adjust the hyperparameters of loss terms during the training process (Xiang et al., 2022), (2) the Bayesian PINNs (B-PINNs) that are specialized to deal with forward and inverse nonlinear problems with noisy data (Yang et al., 2021), (3) Rectified PINNs (RPINNs) that are trained with the gradient information from the numerical solution by the multigrid method and designed for solving stationary PDEs  (Peng et al., 2022), (4) Auxiliary Pinns (A-PINNs) that effectively handle integro-differential equations (Yuan et al., 2022), (5) conservative PINNs (cPINNs) and exetended PINNs (XPINNs) that adopt the domain decomposition technique (Jagtap et al., 2020; Jagtap & Karniadakis, 2020), (6) parrel PINNs that reduces the computational cost of cPINNs and XPINNs (Shukla et al., 2021), (7) gradient-enhanced PINNs (gPINNs) that use the gradient of the PDE loss term with respect to the network inputs  (Yu et al., 2022).
-PINNs primarily employ Xavier initialization for training (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023; Song et al., 2024), but our experimental results indicate that this method limits the performance of larger network sizes.
-Although there have been recent results on initialization methods for PINNs, most of them have relied on transfer learning (Tarbiyati & Nemati Saray, 2023).
-Thus, we propose a weight initialization method that does not require transfer learning and is robust to variations in network size.
+Several initialization methods have been proposed to improve the training of deep ReLU networks (Lu et al., 2019; Bachlechner et al., 2021; Zhao et al., 2022; Lee et al., 2024).
+However, research on weight initialization for tanh-based neural networks remains relatively limited.
+Nevertheless, tanh networks have gained increasing attention in recent years, particularly in applications such as PINNs, where their performance is highly sensitive to initialization randomness.
+PINNs have shown promising results in solving forward, inverse, and multiphysics problems in science and engineering.
+(Mao et al., 2020; Shukla et al., 2020; Lu et al., 2021; Karniadakis et al., 2021; Yin et al., 2021; Bararnia & Esmaeilpour, 2022; Cuomo et al., 2022; Hanna et al., 2022; Hosseini et al., 2023; Wu et al., 2023; Zhu et al., 2024).
+PINNs approximate solutions to partial differential equations (PDEs) using neural networks and are trained by minimizing a loss function, typically formulated as a sum of least-squares terms incorporating PDE residuals, boundary conditions, and initial conditions.
+This loss is commonly optimized using gradient-based methods such as Adam (Kingma & Ba, 2014), L-BFGS (Liu & Nocedal, 1989), or a combination of both.
+Universal approximation theorems (Cybenko, 1989; Hornik et al., 1989; Hornik, 1991; Park et al., 2021; Guliyev & Ismailov, 2018b; Shen et al., 2022; Guliyev & Ismailov, 2018a; Maiorov & Pinkus, 1999; Yarotsky, 2017; Gripenberg, 2003) establish the theoretical capability of neural networks to approximate analytic PDE solutions.
+However, PINNs still face challenges in accuracy, stability, computational complexity, and tuning appropriate hyperparameters of loss terms.
+To address these challenges, various improved variants of PINNs have been proposed: (1) Self-adaptive loss-balanced PINNs (lbPINNs), which automatically adjust the hyperparameters of loss terms during training (Xiang et al., 2022), (2) Bayesian PINNs (B-PINNs), designed to handle forward and inverse nonlinear problems with noisy data (Yang et al., 2021), (3) Rectified PINNs (RPINNs), which incorporate gradient information from numerical solutions obtained via the multigrid method and are specifically designed for solving stationary PDEs (Peng et al., 2022), (4) Auxiliary PINNs (A-PINNs), developed to effectively handle integro-differential equations (Yuan et al., 2022), (5) Conservative PINNs (cPINNs) and Extended PINNs (XPINNs), which employ domain decomposition techniques (Jagtap et al., 2020; Jagtap & Karniadakis, 2020), (6) Parallel PINNs, designed to reduce the computational cost of cPINNs and XPINNs (Shukla et al., 2021), and (7) Gradient-enhanced PINNs (gPINNs), which incorporate the gradient of the PDE loss term with respect to network inputs (Yu et al., 2022).
+While these advancements address various challenges in PINNs, activation functions and their initialization strategies remain crucial for achieving optimal performance.The tanh activation function has been shown to perform well in PINNs (Raissi et al., 2019), with detailed experimental results provided in Appendix C.2.
+Xavier initialization is widely employed as the standard approach for tanh networks in existing studies (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023; Song et al., 2024).
+However, our experimental results indicate that employing Xavier initialization leads to decreased model performance as network size increases.
+Moreover, performance gains from Batch Normalization or Layer Normalization remain limited, and Xavier initialization exhibits sensitivity to the amount of training data, particularly in smaller datasets.
+Although a recent study has proposed an initialization method for PINNs, it depends on transfer learning (Tarbiyati & Nemati Saray, 2023).
+To address these limitations, we propose a weight initialization method that is robust across different network sizes, achieves high data efficiency, and reduces reliance on both transfer learning and normalization techniques.
 (a) Xavier Initialization
 (b) Proposed Initialization
-Figure 1: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using Xavier initialization (Left) and the proposed initialization (Right).
-Experiments were conducted on networks with 10,000 hidden layers, each having the same number of nodes: $16$, $32$, $64$, or $128$.
+Figure 1: The difference between the maximum and minimum activation values at each layer when propagating $3,000$ input samples through a $10,000$-layer tanh FFNN, using Xavier initialization (left) and the proposed initialization (right).
+Experiments were conducted on distinct networks with $10,000$ hidden layers, each having the same number of nodes: $16$, $32$, $64$, or $128$.
 3 Proposed Weight Initialization method
 In this section, we discuss the proposed weight initialization method.
-Subsection 3.1 introduces the theoretical motivation behind the methodology.
-Subsection 3.2 presents how to derive the initial weight matrix that satisfies the conditions outlined in Subsection 3.1.
-Finally, in Subsection 3.3, we suggest the optimal hyperparameter $\sigma_{z}$ in the proposed method.
+Section 3.1 introduces the theoretical motivation behind the method.
+Section 3.2 presents how to derive the initial weight matrix that satisfies the conditions outlined in Section 3.1.
+Finally, in Section 3.3, we suggest the optimal hyperparameter $\sigma_{z}$ in the proposed method.
 3.1 Theoretical motivation
-Experimental results in Figure 1 reveal that when Xavier initialization is employed in FFNNs with tanh activation, the distribution of activation values tends to cluster around zero in deeper layers.
+Experimental results in Figure 1 indicate that when Xavier initialization is employed in FFNNs with tanh activation, the activation values tend to vanish toward zero in deeper layers.
 This vanishing of activation values can hinder the training process due to a discrepancy between the activation values and the desired output.
-However, theoretically preventing this phenomena is not straightforward.
-In this subsection, we gives a theoretical analysis based on a fixed point of $\tanh(ax)$ to bypass the phenomena.
-Before giving the theoretical foundations, consider the basic results for a tanh activation function.
-Recall that $x^{*}$ is a fixed point of a function $f$ if $x^{*}$ belongs to both the domain and the codomain of $f$, and $f(x^{*})=x^{*}$.
-The proofs of Lemma 1 and Lemma 2 are provided in Appendix A.
+However, it is not straightforward to theoretically establish the conditions for preventing this phenomenon.
+In this section, we present a theoretical analysis based on a fixed point of $\tanh(ax)$ to mitigate this issue.
+Before giving the theoretical foundations, consider the basic results for a hyperbolic tangent activation function.
+Recall that if $f:\mathbb{R}\to\mathbb{R}$ is a function, then an element $x^{*}\in\mathbb{R}$ is called a fixed point of $f$ if $f(x^{*})=x^{*}$.
 Lemma 1.
-For a fixed $a>0$ define the function $\phi_{a}:\mathbb{R}\to\mathbb{R}$ given as
+For a fixed $a>0$, define the function $\phi_{a}:\mathbb{R}\to\mathbb{R}$ given as
 $\phi_{a}(x):=\tanh(ax).$
-Then, there exists a fixed point $x^{\ast}$. Furthermore,
+Then, there exists a fixed point $x^{\ast}$ of $\phi_{a}$. Furthermore,
   • $(1)$
-    if $0<a\leq 1$, then $\phi$ has a unique fixed point $x^{\ast}=0$.
+    if $0<a\leq 1$, then $\phi_{a}$ has a unique fixed point $x^{\ast}=0$.
   • $(2)$
-    if $a>1$, then $\phi$ has three distinct fixed points: $x^{\ast}=-\xi_{a}$, 00, $\xi_{a}$ such that $\xi_{a}>0$.
-Remark that the function $\phi_{a}$ can be considered as one-layer tanh FFNN.
+    if $a>1$, then $\phi_{a}$ has three distinct fixed points: $x^{\ast}=-\xi_{a}$, 00, $\xi_{a}$ such that $\xi_{a}>0$.
+Proof.
+The proof is detailed in Appendix A.1. ∎
+Note that $\tanh(x)<x$ for all $x>0$.
+However, according to Lemma 1, for $a>1$ the behavior of $\tanh(ax)$ changes.
+If $x>\xi_{a}$ (resp.
+$x<\xi_{a}$), then $\tanh(ax)<x$ (resp.
+$\tanh(ax)>x$).
+At $x=\xi_{a}$, the equality $\tanh(ax)=x$ holds.
+The following lemma addresses the convergence properties of iteratively applying $\tanh(ax)$ for any $x>0$.
 Lemma 2.
-For a given initial value $x_{0}>0$ define
+For a given initial value $x_{0}>0$, define
 $x_{n+1}=\phi_{a}(x_{n}),\quad n=0,1,2,\ldots.$
@@
     if $a>1$, then $x_{n}\to\xi_{a}$ as $n\to\infty$.
-Note that the parameter $a$ in Lemma 2 does not change across all iterations. In Propositions 3 and Corollary 4, we address cases where the value of $a$ varies with each iteration.
+Proof.
+The proof is detailed in Appendix A.2. ∎
+Note that for $a>1$ and $x_{0}>\xi_{a}>0$, the sequence $\{x_{n}\}$ satisfies $\tanh(ax_{n})>\tanh(ax_{n+1})>\xi_{a}$ for all $n\in\mathbb{N}$.
+Similarly, when $0<x_{0}<\xi_{a}$, the sequence satisfies $\tanh(ax_{n})<\tanh(ax_{n+1})<\xi_{a}$ for all $n\in\mathbb{N}$.
+Given $a>1$ and $x_{0}<0$, the sequence converges to $-\xi_{a}$ as $n\to\infty$ due to the odd symmetry of $\tanh(ax)$.
+According to Lemma 2, when $a>1$, the sequence $\{x_{n}\}$ converges to $\xi_{a}$ or $-\xi_{a}$, respectively, as $n\to\infty$, for an arbitrary initial value $x_{0}>0$ or $x_{0}<0$.
+Note that the parameter $a$ in Lemma 2 does not change across all iterations. Propositions 3 and Corollary 4, consider the case where the value of $a$ varies for each iteration.
 Proposition 3.
@@
 Proof.
-Set $N=\max\{n|a_{n}>1\}$.
-Define the sequences $\{b_{n}\}_{n=1}^{\infty}$ and $\{c_{n}\}_{n=1}^{\infty}$ such that $b_{n}=c_{n}=a_{n}$ for $n\leq N$, with $b_{n}=0$ and $c_{n}=1$ for $n>N$.
-Suppose that $\{\hat{\Phi}_{m}\}_{m=1}^{\infty}$ and $\{\tilde{\Phi}_{m}\}_{m=1}^{\infty}$ are sequences of functions defined as for each $m\in\mathbb{N}$
-$\hat{\Phi}_{m}=\phi_{b_{m}}\circ\phi_{b_{m-1}}\circ\cdots\circ\phi_{b_{1}},\quad\tilde{\Phi}_{m}=\phi_{c_{m}}\circ\phi_{c_{m-1}}\circ\cdots\circ\phi_{c_{1}}.$
-Then, the inequality $\hat{\Phi}_{m}\leq\Phi_{m}\leq\tilde{\Phi}_{m}$ holds for all $m$.
-By Lemma 1, for any $x\geq 0$, we have $\lim_{m\to\infty}\hat{\Phi}_{m}=0$ and $\lim_{m\to\infty}\tilde{\Phi}_{m}=0$.
-Therefore, the Squeeze Theorem guarantees that $\lim_{m\to\infty}\Phi_{m}(x)=0$.
+The proof is detailed in Appendix A.3. ∎
 Corollary 4.
@@
 Then for any $x\in\mathbb{R}\setminus\{0\}$
-$\left|\lim_{m\to\infty}\Phi_{m}(x)\right|\geq\xi_{1+\epsilon}$
+$\lim_{m\to\infty}\left|\Phi_{m}(x)\right|\geq\xi_{1+\epsilon}.$
 Proof.
-Set $N=\max\{n\mid a_{n}<1+\epsilon\}$.
-Define the sequence $\{b_{n}\}_{n=1}^{\infty}$ such that $b_{n}=a_{n}$ for $n\leq N$, and $b_{n}=1+\epsilon$ for $n>N$.
-The remainder of the proof is analogous to the proof of Proposition 3.
-∎
+The proof is detailed in Appendix A.4. ∎
+Based on Proposition 3 and Corollary 4, if there exists a sufficiently large $N\in\mathbb{N}$ such that $a_{n}<1$ (resp.
+$a_{n}>1+\epsilon$) for all $n\geq N$, then for any $x_{0}\in\mathbb{R}\setminus\{0\}$, $\Phi_{m}(x_{0})\to 0$ (resp.
+$|\Phi_{m}(x_{0})|\geq\xi_{1+\epsilon}$) as $m\to\infty$.
+This result implies that if the sequence $\{\Phi_{m}\}_{m=1}^{M}$ is finite, $a_{n}$ for $N\leq n\leq M$, where $N$ is an arbitrarily chosen index close to $M$, significantly influence the values of $\Phi_{M}(x_{0})$.
 3.2 The derivation of the proposed weight initialization method
-To establish the notation, consider a feedforward neural network with $L$ layers.
+Based on the theoretical motivations discussed in the previous section, we propose a weight initialization method that satisfies the following conditions in the initial forward pass:
+  • (i)
+    It prevents activation values from vanishing towards zero in deep layers.
+  • (ii)
+    It ensures that the distribution of activation values in deep layers is approximately normal.
+Notations.
+Consider a feedforward neural network with $L$ layers.
 The network processes $K$ training samples, denoted as pairs $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{K}$, where $\bm{x}_{i}\in\mathbb{R}^{N_{x}}$ is training input and $\bm{y}_{i}\in\mathbb{R}^{N_{y}}$ is its corresponding output.
@@
 where $\mathbf{W}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ is the weight matrix, $\mathbf{b}^{\ell}\in\mathbb{R}^{N_{\ell}}$ is the bias, and $\tanh(\cdot)$ is an element-wise activation hyperbolic tangent function.
+Denote $\mathbf{W}^{\ell}=[w_{ij}^{\ell}]$.
+Signal Propagation Analysis.
 We present a simplified analysis of signal propagation in FFNNs with the tanh activation function.
 For notational convenience, it is assumed that all hidden layers, as well as the input and output layers, have a dimension of $n$, i.e., $N_{\ell}=n$ for all $\ell$.
-Given an arbitrary input vector $\bm{x}=(x_{1},\ldots,x_{n})$, the first layer activation $\bm{x}^{1}=\tanh(\mathbf{W}^{1}\bm{x})$ can be expressed component-wise as:
-$x^{1}_{i}=\tanh\big{(}w^{1}_{i1}x_{1}+\cdots+w^{1}_{in}x_{n}\big{)}=\tanh\bigg{(}\Big{(}w^{1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{1}_{ij}x_{j}}{x_{i}}\Big{)}x_{i}\bigg{)},\,\text{for }i=1,\dots,n.$
-For the $k+1$-th layer, $i=1,\ldots,n$, this expression can be generalized as:
-$x_{i}^{k+1}=\tanh\left(a_{i}^{k+1}x_{i}^{k}\right),\,\text{where}\quad a_{i}^{k+1}=w^{k+1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{k+1}_{ij}x_{j}^{k}}{x_{i}^{k}}.$ (1)
+Given an arbitrary input vector $\bm{x}=(x_{1},\ldots,x_{n})$, the first layer activation $\bm{x}^{1}=\tanh(\mathbf{W}^{1}\bm{x})$ can be expressed component-wise as
+$x^{1}_{i}=\tanh\big{(}w^{1}_{i1}x_{1}+\cdots+w^{1}_{in}x_{n}\big{)}=\tanh\bigg{(}\Big{(}w^{1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{1}_{ij}x_{j}}{x_{i}}\Big{)}x_{i}\bigg{)}\text{ for }i=1,\dots,n.$
+For the ($k+1$)-th layer this expression can be generalized as
+$x_{i}^{k+1}=\tanh\left(a_{i}^{k+1}x_{i}^{k}\right),\,\text{where}\quad a_{i}^{k+1}=w^{k+1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{k+1}_{ij}x_{j}^{k}}{x_{i}^{k}}\quad\text{for }i=1,\ldots,n.$ (1)
+Equation (1) follows the form of $\tanh(ax)$, as discussed in Section 3.2.
 According to Lemma 2, when $a>1$, for an arbitrary initial value $x_{0}>0$ or $x_{0}<0$, the sequence $\{x_{k}\}$ defined by $x_{k+1}=\tanh(ax_{k})$ converges to $\xi_{a}$ or $-\xi_{a}$, respectively, as $k\to\infty$.
-This result indicates that the sequence converges to the fixed point $\xi_{a}$ regardless of the initial value $x_{0}$ and ensures that the activation values do not vanish as network depth increases.
-Furthermore, by Lemma 2, if $a^{k}_{i}\leq 1$ for all $N\leq k\leq L$, then $x_{i}^{L}$ approaches zero.
-Therefore, to ensure that ((((i)))) $a_{i}^{k}$ remains close to 1 and ((((ii)))) $a_{i}^{k}\leq 1$ does not hold for all $N\leq k\leq L$, we design the initial weight matrix as $\mathbf{W}^{\ell}=\mathbf{D}^{\ell}+\mathbf{Z}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$, where $\mathbf{D}^{\ell}_{i,j}=1$ if $i\equiv j\pmod{N_{\ell-1}}$, 0 otherwise, and $\mathbf{Z}^{\ell}$ is a noise matrix drawn from $\mathcal{N}(0,\sigma_{z}^{2})$, where $\sigma_{z}$ is set to $\alpha/\sqrt{N^{\ell-1}}$ and $\alpha=0.085$.
+This result indicates that the sequence converges to the fixed point $\xi_{a}$ regardless of the initial value $x_{0}$.
+From the perspective of signal propagation in tanh-based FFNNs, this ensures that the activation values do not vanish as the network depth increases.
+Furthermore, by Proposition 3, if $a^{k}_{i}\leq 1$ for all $N\leq k\leq L$, where $N$ is an arbitrarily chosen index sufficiently close to $L$, the value of $x_{i}^{L}$ approaches zero.
+Therefore, to satisfy condition ((((i)))), $a_{i}^{k}$ remains close to 1, and the inequality $a_{i}^{k}\leq 1$ does not hold for all $N\leq k\leq L$.
+Proposed Weight Initialization.
+The proposed initial weight matrix is defined as $\mathbf{W}^{\ell}=\mathbf{D}^{\ell}+\mathbf{Z}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$, where $\mathbf{D}^{\ell}_{i,j}=1$ if $i\equiv j\pmod{N_{\ell-1}}$, and 0 otherwise (Examples of $\mathbf{D}^{\ell}$ are provided in Appendix D).
+The noise matrix $\mathbf{Z}^{\ell}$ is drawn from $\mathcal{N}(0,\sigma_{z}^{2})$, where $\sigma_{z}$ is set to $\alpha/\sqrt{N^{\ell-1}}$ with $\alpha=0.085$.
 Then $a_{i}^{k+1}$ follows the distribution:
 $a_{i}^{k+1}\sim\mathcal{N}\Bigg{(}1,\,\sigma_{z}^{2}+\sigma_{z}^{2}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\bigg{(}\frac{x_{j}^{k}}{x_{i}^{k}}\bigg{)}^{2}\Bigg{)}.$ (2)
-According to Equation 2, the mean of $a_{i}^{k+1}$ is $1$, so choosing an appropriate $\sigma_{z}$ satisfies condition ((((i)))).
-For condition ((((ii)))), if $x_{i}^{k}$ becomes small relative to other elements in $\bm{x}^{k}$, the variance of $a_{i}^{k+1}$ increases, as indicated by Equation 2.
-As a result, the probability that the absolute value of $x_{i}^{k+1}$ surpasses that of $x_{i}^{k}$ is higher.
-However, if $\sigma_{z}$ is too small, the increase in the variance of $a_{i}^{k+1}$ becomes limited.
-Therefore, choosing an appropriate $\sigma_{z}$ is crucial.
+According to Equation (2), $a_{i}^{k+1}$ follows a Gaussian distribution with a mean of $1$.
+Additionally, if $x_{i}^{k}$ becomes small relative to other elements in $\bm{x}^{k}$, the variance of the distribution in (2) increases.
+Consequently, the probability that the absolute value of $x_{i}^{k+1}$ exceeds that of $x_{i}^{k}$ becomes higher.
+Figure 1 (b) shows that activation values maintain consistent scales in deeper layers.
 (a) Same dimension
 (b) Varying dimensions
-Figure 2: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using the proposed initialization with $\alpha$ set to $0.04,0.085,0.15,$ and $0.5$.
+Figure 2: The difference between the maximum and minimum activation values at each layer when propagating $3,000$ input samples through a $10,000$-layer tanh FFNN, using the proposed initialization with $\alpha$ set to $0.04,0.085,0.15,$ and $0.5$.
 Network with $10,000$ hidden layers, each with 32 nodes (left), and a network with alternating hidden layers of $64$ and $32$ nodes (right).
 3.3 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
-In this subsection, we discuss how $\sigma_{z}$ impacts the scale of the activation values.
-Equation 2 indicates that $a_{i}^{k}$ follows a normal distribution, with variance depending on $\sigma_{z}$.
+In this section, we determine the appropriate value of $\alpha$ in $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ that satisfies condition ((((ii)))).
+Condition ((((ii)))) is motivated by normalization methods (Ioffe, 2015; Ba, 2016).
 Firstly, we experimentally investigated the impact of $\sigma_{z}$ on the scale of the activation values.
-As demonstrated in Figure 2, increasing $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ causes the activation values in any layer to be distributed over a broader range.
-However, setting $\sigma_{z}$ to a large value can lead to saturation, where most activations converge towards $-1$ and $1$.
-If $\sigma_{z}$ is too large, the probability that $a_{i}^{(k)}$ takes values far from $1$ (e.g., -10, 5, etc.) increases.
-This, in turn, increases the value of $1+\epsilon$ mentioned in Corollary 4, potentially bounding the activation values in sufficiently deep layers by $\xi_{1+\epsilon}$.
-Consequently, the activation values in deeper layers become less likely to approach zero and tend to saturate toward specific values.
-On the other hand, if $\sigma_{z}$ is too small, as mentioned in Subsection 3.2, the variance of $a_{i}^{k}$ becomes restricted.
-This is demonstrated experimentally in Figure 2, when $\alpha=0.00001$.
-For this reason, we experimentally found an optimal $\sigma_{z}=\alpha/\sqrt{N^{\ell-1}}$, with $\alpha=0.085$, that is neither too large nor too small.
-Results from experiments solving the Burgers’ equation using PINNs with varying $\sigma_{z}$ are presented in Appendix B.1.
-Table 1: Validation accuracy and loss are presented for FFNNs with varying numbers of nodes $(2,8,32,128)$, each with $20$ hidden layers using the tanh activation function.
-All models were trained for $20$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented.
-When comparing different initialization methods under the same experimental settings, the better-performing method is highlighted in bold.
-Underlined values indicate the highest accuracy when only the number of nodes is varied.
-2 8 32 128
-MNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
-Xavier 49.78 1.632 68 0.958 91.67 0.277 95.45 0.154
-Proposed 62.82 1.185 77.95 0.706 92.51 0.255 96.12 0.134
-FMNIST Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
-Xavier 42.89 1.559 68.55 0.890 81.03 0.533 86.2 0.389
-Proposed 51.65 1.324 71.31 0.777 83.06 0.475 87.12 0.359
-CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
-Xavier 32.82 1.921 43.51 1.608 48.62 1.473 47.58 1.510
-Proposed 38.16 1.780 47.04 1.505 48.80 1.463 48.51 1.471
-Table 2: Validation accuracy and loss are presented for FFNNs with varying numbers of layers $(10,50,100)$, each with $64$ number of nodes using the tanh activation function.
-All models were trained for $40$ epochs, and the highest average accuracy and lowest average loss, computed from $10$ runs, are presented.
-10 50 100
-MNIST Accuracy Loss Accuracy Loss Accuracy Loss
-Xavier 96.55 0.112 96.57 0.123 94.08 0.194
-Proposed 97.04 0.102 96.72 0.109 96.06 0.132
-FMNIST Accuracy Loss Accuracy Loss Accuracy Loss
-Xavier 88.73 0.319 87.72 0.344 83.41 0.463
-Proposed 89.42 0.305 88.51 0.324 86.01 0.382
-CIFAR10 Accuracy Loss Accuracy Loss Accuracy Loss
-Xavier 48.39 1.468 47.87 1.474 46.71 1.503
-Proposed 48.41 1.458 48.71 1.461 48.96 1.437
+As shown in Figure 2, increasing $\sigma_{z}=\alpha/\sqrt{N_{\ell-1}}$ broadens the activation range in each layer, while decreasing $\sigma_{z}$ narrows it.
+When $\sigma_{z}$ is Large.
+Setting $\sigma_{z}$ to a large value can lead to saturation.
+If $\sigma_{z}$ is too large, Equation (2) implies that the likelihood of $a_{i}^{k}$ deviating significantly from 1 increases.
+This increases the likelihood of activation values being bounded by $\xi_{1+\epsilon}$ in sufficiently deep layers, as stated in Corollary 4.
+Consequently, in deeper layers, activation values are less likely to approach zero and tend to saturate toward specific values.
+Please refer to the Figure 3 for the cases where $\sigma_{z}=0.3$ and $3$.
+Figure 3: The activation values in the 1000th layer, with $32$ nodes per hidden layer, were analyzed using the proposed weight initialization method with $\sigma_{z}$ values of $0.0003$, $0.015$, $0.3$, and $3$.
+The analysis was conducted on $3,000$ input samples uniformly distributed within the range $[-1,1]$.
+When $\sigma_{z}$ is Small.
+If $\sigma_{z}$ is too small, Equation (2) implies that the distribution of $a_{i}^{k}$ has a standard deviation close to zero.
+Consequently, $x_{i}^{k+1}$ can be approximated as the result of applying $\tanh(x)$ to $x_{i}>0$ repeatedly for a finite number of iterations, $k$.
+Since $\tanh^{\prime}(x)$ decreases for $x\geq 0$, the values resulting from finite iterations eventually saturate.
+Plese refer to the Figure 3 when $\sigma_{z}=0.0003$.
+For these reasons, we experimentally determined an optimal $\sigma_{z}$ that avoids being excessively large or small.
+As shown in Figure 3, $\sigma_{z}=0.015$ maintains an approximately normal activation distribution without collapse.
+Additional experimental results are provided in Appendix A.5.
+Considering the number of hidden layer nodes, we set $\sigma_{z}=\alpha/\sqrt{N^{\ell-1}}$ with $\alpha=0.085$.
+Experimental results for solving the Burgers’ equation using PINNs with varying $\sigma_{z}$ are provided in Appendix C.3.
+Table 1: Validation accuracy and loss are presented for FFNNs with varying numbers of nodes $(2,8,32,128,512)$, each with $20$ hidden layers using tanh activation function.
+All models were trained for $20$ epochs, and the highest average accuracy and lowest average loss, computed over $10$ runs, are presented.
+The better-performing method is highlighted in bold when comparing different initialization methods under the same experimental settings.
+Dataset Method 2 Nodes 8 Nodes 32 Nodes 128 Nodes 512 Nodes
+Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
+MNIST Xavier 49.78 1.632 68 0.958 91.67 0.277 95.45 0.154 97.35 0.087
+Proposed 62.82 1.185 77.95 0.706 92.51 0.255 96.12 0.134 97.96 0.067
+FMNIST Xavier 42.89 1.559 68.55 0.890 81.03 0.533 86.20 0.389 88.28 0.331
+Proposed 51.65 1.324 71.31 0.777 83.06 0.475 87.12 0.359 88.59 0.323
+CIFAR-10 Xavier 32.82 1.921 43.51 1.608 48.62 1.473 47.58 1.510 51.71 1.369
+Proposed 38.16 1.780 47.04 1.505 48.80 1.463 48.51 1.471 52.21 1.359
+CIFAR-100 Xavier 10.87 4.065 18.53 3.619 23.71 3.301 23.83 3.324 17.72 3.672
+Proposed 15.22 3.818 23.07 3.350 24.93 3.237 24.91 3.240 22.80 3.435
+Table 2: Validation accuracy and loss are presented for FFNNs with varying numbers of layers $(3,10,50,100)$, each with $64$ number of nodes using the tanh activation function.
+All models were trained for $40$ epochs, and the highest average accuracy and lowest average loss, computed over $10$ runs, are presented.
+Dataset Method 3 Layers 10 Layers 50 Layers 100 Layers
+Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
+MNIST Xavier 95.98 0.130 96.55 0.112 96.57 0.123 94.08 0.194
+Proposed 96.32 0.123 97.04 0.102 96.72 0.109 96.06 0.132
+FMNIST Xavier 85.91 0.401 88.73 0.319 87.72 0.344 83.41 0.463
+Proposed 86.51 0.379 89.42 0.305 88.51 0.324 86.01 0.382
+CIFAR-10 Xavier 42.91 1.643 48.39 1.468 47.87 1.474 46.71 1.503
+Proposed 45.05 1.588 48.41 1.458 48.71 1.461 48.96 1.437
+CIFAR-100 Xavier 19.10 3.628 22.73 3.400 24.27 3.283 20.32 3.515
+Proposed 19.30 3.609 23.83 3.309 25.07 3.190 24.41 3.234
 4 Experiments
-In this section, we conduct a series of experiments to validate the proposed weight initialization method.
-In Subsection 4.1, we evaluate the performance of an FFNN with the tanh activation function on benchmark datasets.
-In Subsection 4.2, we solve the Burgers’ equation and Allen-Cahn equation using Physics-Informed Neural Networks.
-Both experiments are conducted across various network sizes to verify whether the proposed method consistently performs well, independent of network depth and width.
-The experiments were conducted in TensorFlow without skip connections, normalization layers, and learning rate decay in any of the experiments.
-4.1 Width Independence in Classification Task
-To evaluate the effectiveness of the proposed weight initialization method, we conduct experiments on the MNIST, Fashion MNIST, and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, utilizing the Adam optimizer.
-All experiments are conducted with a batch size of 64 and a learning rate of 0.0001.
-Fifteen percent of the total dataset is allocated for validation.
-We apply the proposed weight initialization method to evaluate its effectiveness in training tanh FFNNs, emphasizing its robustness to variations in network width.
-Four tanh FFNNs are created, each with 20 hidden layers, and with 2, 8, 32, and 128 nodes per hidden layer, respectively.
-In Table 1, for both the MNIST and Fashion MNIST datasets, the network with 128 nodes achieves the highest accuracy and lowest loss when our proposed method is employed.
-However, for the CIFAR-10 dataset, the network with 32 nodes yields the highest accuracy and lowest loss when employing the proposed method.
-In summary, our proposed method demonstrates robustness regardless of the number of nodes in tanh FFNNs.
-We provide more detailed experimental results in Appendix B.2.
-4.2 Depth Independence in Classification Task
-It is well known that the expressivity of neural networks generally increases exponentially with depth, enabling strong generalization performance (Poole et al., 2016; Raghu et al., 2017).
-Therefore, we employ the proposed weight initialization method to investigate its effectiveness in training deep FFNNs with the tanh activation function, emphasizing its robustness to variations in network depth.
-We create three tanh FFNNs, each with 64 nodes in all hidden layers, but with 10, 50, and 100 hidden layers, respectively.
+In this section, we present a series of experiments to evaluate the proposed weight initialization method.
+In Section 4.1, we evaluate the performance of tanh FFNNs on benchmark classification datasets.
+In Section 4.2, we apply Physics-Informed Neural Networks to solve PDEs.
+Both experiments assess the proposed method’s robustness to network size and data efficiency.
+4.1 Classification Task
+Experimental Setting.
+To evaluate the effectiveness of the proposed weight initialization method, we conduct experiments on the MNIST, Fashion MNIST (FMNIST), CIFAR-10, and CIFAR-100 (Krizhevsky & Hinton, 2009) datasets with the Adam optimizer.
+All experiments are performed with a batch size of $64$ and a learning rate of $0.0001$.
+Fifteen percent of the dataset is allocated for validation.
+The experiments are implemented in TensorFlow without skip connections and learning rate decay in any of the experiments.
+Width Independence in Classification Task.
+We evaluate the proposed weight initialization method in training tanh FFNNs, focusing on its robustness across different network widths.
+Five tanh FFNNs are designed, each with 20 hidden layers, and with $2,8,32$, $128$, and $512$ nodes per hidden layer, respectively.
+In Table 1, for the MNIST, Fashion MNIST and CIFAR-10 datasets, the network with 512 nodes achieves the highest accuracy and lowest loss when the proposed method is employed.
+However, for the CIFAR-100 dataset, the network with 32 nodes yields the highest accuracy and lowest loss when employing the proposed method.
+In summary, the proposed method demonstrates robustness across different network widths in tanh FFNNs.
+Detailed experimental results are provided in Appendix B.1.
+Depth Independence in Classification Task.
+The expressivity of neural networks is known to increase exponentially with depth, leading to improved generalization performance (Poole et al., 2016; Raghu et al., 2017).
+To evaluate the robustness of the proposed weight initialization method across different network depths, we conduct experiments on deep FFNNs with tanh activation functions.
+Specifically, we construct four tanh FFNNs, each with $64$ nodes per hidden layer and $3$, $10$, $50$, and $100$ hidden layers.
+respectively.
 In Table 2, for both the MNIST and Fashion MNIST datasets, the network with 10 hidden layers achieves the highest accuracy and lowest loss when our proposed method is employed.
-Both initialization methods perform best in networks with the fewest layers, with performance degrading as the depth increases.
-However, for the CIFAR-10 dataset, we observe that the performance of the proposed method improves as the number of layers increases.
-Furthermore, we conduct experiments with varying hidden layer dimensions, as shown in Figure 3.
-The network consists of 60 hidden layers, where the number of nodes alternates between 32 and 16 in each layer.
-We demonstrate superior performance in terms of both loss and accuracy across all epochs on the MNIST and CIFAR-10 datasets.
+Both initialization methods exhibit lower performance in networks with $3$ layers compared to those with more layers.
+Moreover, for more complex datasets such as CIFAR-10 and CIFAR-100, the proposed method demonstrated improved performance when training deeper networks.
+(a) MNIST
+(b) Fashion MNIST
+(c) CIFAR10
+(d) CIFAR100
+Figure 4: Validation accuracy for a tanh FFNN with 50 hidden layers (32 nodes each).
+Xavier + BN and Xavier + LN represent Xavier initialization with Batch Normalization or Layer Normalization applied every 5 layers, respectively.
+Normalization Methods.
+Xavier initialization is known to cause vanishing gradients and activation problems in deeper networks.
+These issues can be mitigated by applying Batch Normalization (BN) or Layer Normalization (LN) in the network.
+Therefore, we compare the proposed method with Xavier, Xavier with BN, and Xavier with LN.
+To evaluate the effectiveness of normalization, we conducted experiments using a sufficiently deep neural network with $50$ hidden layers.
+As shown in Figure 4, for datasets with relatively fewer features, such as MNIST and FMNIST, Xavier with normalization converges faster than Xavier.
+However, for feature-rich datasets such as CIFAR-10 and CIFAR-100, the accuracy of Xavier with normalization is lower than that of Xavier.
+Normalization typically incurs a $30\%$ computational overhead, and additional hyperparameter tuning is required to determine which layers should apply normalization.
+In contrast, the proposed method achieves the best performance across all datasets without requiring normalization.
+Data Efficiency in Classification Task.
+Based on the results in Table 2, we evaluated data efficiency on a network with $50$ hidden layers, each containing $64$ nodes, where Xavier exhibited strong performance.
+As shown in Table 3, the highest average accuracy and lowest average loss over $5$ runs after $100$ epochs are presented for datasets containing $10$, $20$, $30$, $50$, and $100$ samples.
+The proposed method achieved the best performance across all sample sizes.
+Non-uniform Hidden Layer Dimensions.
+We evaluate the performance of the proposed initialization in networks where hidden layer dimensions are not uniform.
+As shown in Figure 5, the network consists of 60 hidden layers, where the number of nodes alternates between 32 and 16 in each layer.
+The proposed method demonstrates improved performance in terms of both loss and accuracy across all epochs on the MNIST and CIFAR-10 datasets.
+Additionally, Appendix B.2 presents experiments on networks with larger variations in the number of nodes.
+Motivated by these results, Appendix B.2 further explores autoencoders with significant differences in hidden layer dimensions.
 (a) MNIST Accuracy
@@
 The model was trained for 20 epochs on the MNIST and CIFAR-10 datasets.
-4.3 Network size independence in pinn
-Xavier initialization is the primary method used for training PINNs  (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023).
-In this section, we experimentally demonstrate that the method’s training performance is highly dependent on randomness and network size.
-Additionally, empirical results are provided demonstrating that the proposed method is more robust to variations in network size.
+Table 3: Validation accuracy and loss for a 10-layer FFNN (64 nodes per layer) trained on datasets containing 10, 20, 30, 50, and 100 samples.
+Results show the highest average accuracy and lowest average loss over 5 runs after 100 epochs.
+Dataset Method 10 20 30 50 100
+Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss Accuracy Loss
+MNIST Xavier 31.13 2.281 35.03 2.078 45.05 1.771 58.45 1.227 64.02 1.139
+Xavier + BN 22.46 2.267 33.73 2.053 37.13 2.042 39.78 1.944 57.51 1.464
+Xavier + LN 28.52 2.411 41.54 1.796 41.94 1.886 54.97 1.362 65.11 1.093
+Proposed 37.32 2.204 46.79 1.656 48.60 1.645 61.54 1.131 68.44 1.043
+FMNIST Xavier 36.16 2.320 41.69 1.814 53.86 1.459 64.53 1.140 63.58 1.048
+Xavier + BN 35.44 2.136 38.58 1.925 40.16 1.819 53.93 1.728 59.78 1.237
+Xavier + LN 34.94 2.362 37.90 1.793 53.27 1.470 59.50 1.198 62.01 1.073
+Proposed 37.31 2.217 49.25 1.651 55.19 1.372 66.14 1.057 67.58 0.914
+4.2 Physics-Informed Neural Networks
+Xavier initialization is the most commonly employed method for training PINNs (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023).
+In this section, we experimentally demonstrate that the proposed method is more robust across different network sizes and achieves higher data efficiency compared to Xavier initialization with or without normalization methods.
+Experimental Setting.
 All experiments on Physics-Informed Neural Networks (PINNs) use full-batch training with a learning rate of 0.001.
-In this section, we solve the Allen-Cahn and Burgers’ equations using a tanh FFNN-based PINN with 20,000 collocation points.
-For the Allen-Cahn equation, the diffusion coefficient is set to $d=0.01$.
-The initial condition is defined as $u(x,0)=x^{2}\cos(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=-1$ and $u(1,t)=-1$, applied over the time interval $t\in[0,1]$.
-Similarly, for the Burgers’ equation, a viscosity coefficient of $\nu=0.01$ is employed.
-The initial condition is given by $u(x,0)=-\sin(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=0$ and $u(1,t)=0$ imposed for $t\in[0,1]$.
-The Allen-Cahn equation is expressed as:
-$\frac{\partial u}{\partial t}-d\frac{\partial^{2}u}{\partial x^{2}}=-\frac{u^{3}+u}{d}$
-where $u(x,t)$ represents the solution, $d$ is the diffusion coefficient, and the nonlinear term $u^{3}-u$ models the phase separation dynamics.
-The Burgers’ equation is given by:
-$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}}$
-where $u(x,t)$ is the velocity field, and $\nu$ is the viscosity coefficient.
-Eight tanh FFNNs are created, each with 16 nodes in all hidden layers, but with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively.
-As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 20.
-However, as the depth increases, the loss gradually rises.
-In contrast, the proposed method achieves the lowest loss at a depth of 50 and maintains a loss of 0.00057 even at a depth of 80 layers.
-For the Burgers’ equation, the proposed method achieves the lowest loss at a depth of 60, while at the same depth, Xavier initialization results in a loss that is an order of magnitude higher (approximately $10^{2}$ difference).
-Next, we double the number of nodes to observe the impact of node size on the loss.
-Eight new tanh FFNNs are created, each with 32 nodes in all hidden layers, and with 5, 10, 20, 30, 40, 50, 60, and 80 hidden layers, respectively.
-As shown in Table 3, for the Allen-Cahn equation, Xavier initialization achieves the lowest loss at a network depth of 30.
-However, as the depth increases, the model becomes untrainable, with a loss of 0.694 at a depth of 80.
-In contrast, the proposed method achieves the lowest loss at a depth of 40 and maintains a loss of 0.00059 even at a depth of 80 layers.
-For the Burgers’ equation, both methods show similar loss values up to a depth of 30.
-Beyond a depth of 40, however, the loss steadily increases with the Xavier method, while the proposed method records the lowest loss at a depth of 50.
+In this section, we solve the Allen-Cahn, Burgers, Diffusion, and Poisson equations using a tanh FFNN-based PINN with 20,000 collocation points.
+Details on the PDEs are provided in Appendix C.1.
+Network Size Independence in PINNs.
+We construct eight tanh FFNNs, each with $16$ nodes per hidden layer and $5,10,20,30,40,50,60,$ or $80$ hidden layers.
+As shown in Table 4, for the Allen-Cahn and Burgers’ equations, Xavier+BN and Xavier+LN achieve the lowest loss at a network depth of 30.
+However, their loss gradually increases as depth grows.
+In contrast, the proposed method achieves the lowest loss at depths of 50 and 60, respectively, maintaining strong learning performance even in deeper networks.
+For the Diffusion and Poisson equations, Xavier+LN achieves the lowest loss at depths of 5 and 10, respectively.
+While all methods exhibit increasing loss as network depth increases, the proposed method consistently maintains lower loss in deeper networks.
+Similar trends are observed with 32 nodes.
+Across all tested network sizes and PDEs, the proposed method consistently achieves the lowest loss.
+The proposed method eliminates the computational overhead and hyperparameter tuning required for normalization methods.
 Table 4: A PINN loss is presented for FFNNs with varying numbers of layers $(5,10,20,30,40,50,60,80)$ using the tanh activation function.
@@
 All models were trained for $300$ iterations using Adam and $300$ iterations using L-BFGS.
-The median PINN loss from the final iteration for the Burgers and Allen–Cahn equations, computed over 5 runs, is presented.
+The median PINN loss at the final iteration for the Burgers, Allen–Cahn, Diffusion, and Poisson equations, computed over 5 runs, is presented.
 Allen-Cahn (16 Nodes) 5 10 20 30 40 50 60 80
 Xavier 9.58e-04 8.16e-04 7.61e-04 1.06e-03 1.1e-03 1.24e-03 3.55e-03 1.81e-03
+Xavier + BN 1.42e-03 8.17e-04 8.56e-04 7.07e-04 7.77e-04 8.87e-04 9.11e-04 2.15e-03
+Xavier + LN 6.29e-01 1.77e-03 6.98e-04 1.27e-03 1.82e-03 6.65e-01 3.29e-01 5.86e-01
 Proposed 9.21e-04 7.29e-04 5.76e-04 5.29e-04 5.37e-04 4.03e-04 4.73e-04 5.77e-04
@@
 Xavier 6.97e-03 1.11e-02 7.9e-03 9.71e-03 2.45e-02 2.65e-02 6.5e-02 5.71e-02
+Xavier + BN 8.07e-03 7.72e-03 6.24e-03 1.70e-02 1.50e-02 1.85e-02 2.91e-02 6.84e-02
+Xavier + LN 3.89e-02 1.88e-02 9.48e-03 9.28e-03 2.46e-02 3.30e-02 6.91e-02 4.42e-02
 Proposed 6.19e-03 5.08e-03 5.28e-03 9.31e-04 3.56e-03 8.27e-04 3.43e-04 2.05e-03
+Diffusion (16 Nodes) 5 10 20 30 40 50 60 80
+Xavier 2.52e-03 4.82e-03 9.69e-03 1.33e-02 2.08e-02 1.50e-02 2.92e-02 7.24e-02
+Xavier + BN 2.89e-03 5.77e-03 1.05e-02 9.65e-03 2.76e-02 1.07e-02 9.07e-03 1.43e-02
+Xavier + LN 1.72e-03 6.10e-03 8.04e-03 9.48e-03 2.14e-02 7.59e-03 2.05e-02 2.21e-02
+Proposed 9.14e-04 2.59e-03 2.40e-03 1.01e-03 1.97e-03 1.21e-03 1.12e-03 1.91e-03
+Poisson (16 Nodes) 5 10 20 30 40 50 60 80
+Xavier 1.52e-02 2.87e-02 1.28e-01 9.82e-02 1.15e-01 1.37e-01 1.82e-01 2.55e-01
+Xavier + BN 1.62e-02 2.02e-02 8.72e-02 1.12e-01 2.45e-01 9.85e-02 1.00e-01 1.34e-01
+Xavier + LN 5.39e-01 4.40e-02 1.34e-01 3.91 2.52e+02 2.58 9.79e+02 N/A
+Proposed 1.37e-02 1.70e-02 4.62e-02 2.43e-02 3.75e-02 4.03e-02 6.07e-02 6.01e-02
 Allen-Cahn (32 Nodes) 5 10 20 30 40 50 60 80
 Xavier 3.13e-01 5.03e-02 3.64e-03 2.37e-03 4.03e-03 5.27e-03 1.73e-02 6.94e-01
+Xavier + BN 4.05e-01 8.85e-04 8.41e-04 7.82e-04 9.97e-04 6.80e-04 9.34e-04 6.94e-01
+Xavier + LN 3.31e-01 2.10e-03 5.99e-04 6.71e-04 1.49e-03 1.29e-03 3.31e-02 6.93e-01
 Proposed 1.04e-03 6.92e-04 5.34e-04 4.26e-04 3.31e-04 3.52e-04 3.85e-04 5.96e-04
@@
 Xavier 1.12e-02 3.53e-03 2.72e-03 1.81e-03 7.60e-03 8.56e-03 9.86e-03 1.66e-01
+Xavier + BN 5.88e-03 1.04e-03 1.79e-03 2.80e-03 5.95e-03 3.66e-02 6.60e-02 1.66e-01
+Xavier + LN 4.31e-02 1.21e-02 1.88e-03 7.22e-03 5.54e-03 8.46e-03 9.04e-03 4.86e-02
 Proposed 4.14e-03 4.11e-03 1.58e-03 1.29e-03 7.96e-04 5.85e-04 9.80e-04 1.47e-03
+Diffusion (32 Nodes) 5 10 20 30 40 50 60 80
+Xavier 1.69e-03 6.85e-03 7.63e-03 4.50e-03 8.98e-03 5.67e-03 6.33e-01 1.59
+Xavier + BN 1.68e-03 2.66e-03 1.08e-02 6.00e-03 8.58e-03 6.60e-03 5.66e-02 1.69e+02
+Xavier + LN 8.16e-04 2.85e-03 8.46e-03 4.57e-03 9.40e-03 1.04e-02 2.42e-01 1.67e+02
+Proposed 2.89e-04 8.03e-04 5.25e-04 5.07e-04 5.33e-04 6.17e-04 9.80e-04 1.53e-03
+Poisson (32 Nodes) 5 10 20 30 40 50 60 80
+Xavier 1.09e-02 1.33e-02 3.13e-02 7.69e-02 6.72e-02 8.90e-02 9.68e+02 1.46e+02
+Xavier + BN 1.14e-02 1.47e-02 2.68e-02 3.55e-02 8.25e-02 8.97e-02 4.50e-02 7.75e-01
+Xavier + LN 2.36e-02 2.18e-02 3.07e-02 3.85e-01 1.40 4.69 2.60 6.14
+Proposed 9.63e-03 8.29e-03 1.41e-02 1.88e-02 1.65e-02 1.85e-02 1.73e-02 3.59e-02
+Data Efficiency in PINNs.
+Based on the results in Table 4, we evaluate data efficiency on a network with $30$ hidden layers, each containing $32$ nodes, where Xavier initialization achieved the lowest PINN loss.
+As shown in Figure 6, for the Burgers equation, the Mean Absolute Error (MAE) of the proposed initialization differs significantly from that of Xavier initialization across varying numbers of collocation points.
+In contrast, for the Diffusion equation, the difference in MAE between the two methods becomes more pronounced when the number of collocation points exceeds $20,000$.
+Additionally, Figure 7 illustrates that increasing the number of collocation points enables PINNs with the proposed initialization to predict solutions with lower absolute error.
+For detailed experiments on the Burgers equation, please refer to Appendix C.4.
+(a) Burgers Equation
+(b) Diffusion Equation
+Figure 6: Mean absolute error between the exact solution and PINN-predicted solution with varying numbers of collocation points.
+The FFNN has $30$ hidden layers ($32$ nodes each) and is trained for $300$ iterations using Adam followed by $300$ iterations using L-BFGS.
+The results are averaged over $5$ experiments.
+Figure 7: Absolute error between the exact solution and the PINN-predicted solution for the Diffusion equation with varying numbers of collocation points $(3000,10000,20000,50000)$ using (upper row) Xavier and (lower row) the proposed initialization.
+The FFNN has 30 hidden layers ($32$ nodes each) and is trained for $300$ iterations using Adam followed by $300$ iterations using L-BFGS.
+The color bar ranges from 00 to $0.05$, with values outside this range shown in white.
 5 Conclusion
-In this paper, we have introduced a novel weight initialization method for tanh FFNNs, grounded in the theoretical analysis of fixed points of the $\tanh(ax)$ function.
-Through our fixed-point analysis, we established conditions under which the vanishing or exploding of activation values can be prevented, even as the depth of the network increases.
-Our proposed method exhibits strong robustness to variations in network size, as demonstrated across a variety of FFNN configurations and benchmark datasets, including MNIST, Fashion MNIST, and CIFAR-10.
-In contrast to Xavier initialization, which struggles to maintain stable performance as network depth increases, the proposed method consistently achieves superior results by preserving activation values.
-Furthermore, we explored the impact of the initialization hyperparameter $\sigma_{z}$ on the distribution of activation values.
-We demonstrated both theoretically and experimentally that the choice of $\sigma_{z}$ plays a significant role in maintaining the proper range of activations, balancing between vanishing and saturation.
-In the context of PINNs, the proposed initialization method shows improved performance in solving PDEs such as the Burgers’ equation and the Allen-Cahn equation.
-By maintaining a stable loss function and achieving faster convergence compared to Xavier initialization, our method demonstrates its practical utility in training networks for physical systems.
-A key advantage of the proposed method lies in its robustness to network depth and width, significantly reducing the need for extensive hyperparameter tuning.
-By maintaining stable performance across varying network configurations, our approach helps to minimize the time and effort spent on searching for optimal network architectures, allowing researchers to focus on model design and other aspects of the training process.
-This makes the proposed method particularly valuable in large-scale and resource-constrained applications where efficient training is critical.
-Appendix A Proofs of the theoretical results
+In this study, we proposed a novel weight initialization method for tanh neural networks, baesed on a theoretical analysis of fixed points of $\tanh(ax)$ function.
+The proposed method is experimentally demonstrated to achieve robustness to variations in network size without normalization methods and to exhibit improved data efficiency.
+Therefore, the proposed weight initialization method reduces the time and effort required for training neural networks.
+Acknowledgments
+This work of Hyunwoo Lee and Hayoung Choi was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+2022R1A5A1033624 and RS-2024-00342939).
+The work of Hyunju Kim was supported by the funds of the Open R&\&&D program of Korea Electric Power Corporation.(No-R23XO03).
+Appendix A Analysis of Signal Propagation
+In this section, we provide proofs for the statements presented in Section 3.1.
+Each proof follows from the fundamental properties of $\tanh(ax)$ and analytical derivations.
+An empirical analysis is performed on the activation distribution for normally and beta-distributed input data.
 A.1 Proof of Lemma 1
 Proof.
-We define $g(x)=\tanh(ax)-x$.
-Since $g(x)$ is continuous, and $g(-M)>0$, $g(M)<0$ for a large real number $M\in\mathbb{R}^{+}$, the Intermediate Value Theorem guarantees the existence of a point $x$ such that $g(x)=0$.
+Define $g(x)=\tanh(ax)-x$. Since $g(x)$ is continuous, and $g(-M)>0$, $g(M)<0$ for a large real number $M$, the Intermediate Value Theorem guarantees the existence of a point $x$ such that $g(x)=0$.
 First, consider the case $0<a\leq 1$.
 Since $0<a\leq 1$, the derivative $g^{\prime}(x)=a\cdot\text{sech}^{2}(ax)-1$ satisfies $-1\leq g^{\prime}(x)\leq a-1<0$ for all $x$.
-Hence, $g(x)$ is strictly decreasing and therefore $g(x)$ has the unique root.
+Hence, $g(x)$ is strictly decreasing and therefore $g(x)$ has a unique root.
 At $x=0$, $\phi(0)=\tanh(a\cdot 0)=0$.
 Hence, $x=0$ is the unique fixed point.
-Let us consider the case $a>1$. For $0<x\ll 1$, $\tanh(ax)-x\approx(a-1)x$. Since $a>1,\tanh(ax)-x>0$. On the other hand, since $|\tanh(ax)|<1$ for all $x$,
+We consider the case $a>1$. For $0<x\ll 1$, $\tanh(ax)-x\approx(a-1)x$. Since $a>1,\tanh(ax)-x>0$. On the other hand, since $|\tanh(ax)|<1$ for all $x$,
 $\lim_{x\to\infty}[-1-x]\leq\lim_{x\to\infty}[\tanh(ax)-x]\leq\lim_{x\to\infty}[1-x].$
@@
 (2) Let $x_{0}<\xi_{a}$.
-Since $\phi^{\prime}(x)$ decreasing for $x\geq 0$, with $\phi^{\prime}(0)>1$ and $\xi_{a}$ is the unique fixed point for $x>0$, it holds that $x_{n}<x_{n+1}<\xi_{a}$ for all $n\in\mathbb{N}$.
+Since $\phi^{\prime}(x)$ is decreasing for $x\geq 0$, with $\phi^{\prime}(0)>1$ and $\xi_{a}$ is the unique fixed point for $x>0$, it holds that $x_{n}<x_{n+1}<\xi_{a}$ for all $n\in\mathbb{N}$.
 Thus, by the monotone convergence theorem, the sequence converges to the fixed point $\xi_{a}$.
 The proof is similar when $x_{0}>\xi_{a}$.
+By the monotone convergence theorem, the sequence also converges to the fixed point $\xi_{a}$.
 ∎
-Appendix B Additional experimental results
-B.1 Preventing activation saturation via appropriate $\sigma_{z}$ tuning
+A.3 Proof of Proposition 3
+Proof.
+Set $N=\max\{n|a_{n}>1\}$.
+Define the sequences $\{b_{n}\}_{n=1}^{\infty}$ and $\{c_{n}\}_{n=1}^{\infty}$ such that $b_{n}=c_{n}=a_{n}$ for $n\leq N$, with $b_{n}=0$ and $c_{n}=1$ for $n>N$.
+Suppose that $\{\hat{\Phi}_{m}\}_{m=1}^{\infty}$ and $\{\tilde{\Phi}_{m}\}_{m=1}^{\infty}$ are sequences of functions defined as for each $m\in\mathbb{N}$
+$\hat{\Phi}_{m}=\phi_{b_{m}}\circ\phi_{b_{m-1}}\circ\cdots\circ\phi_{b_{1}},\quad\tilde{\Phi}_{m}=\phi_{c_{m}}\circ\phi_{c_{m-1}}\circ\cdots\circ\phi_{c_{1}}.$
+Then the inequality $\hat{\Phi}_{m}\leq\Phi_{m}\leq\tilde{\Phi}_{m}$ holds for all $m$.
+By Lemma 1, for any $x\geq 0$, we have $\lim_{m\to\infty}\hat{\Phi}_{m}=0$ and $\lim_{m\to\infty}\tilde{\Phi}_{m}=0$.
+Therefore, the Squeeze Theorem guarantees that $\lim_{m\to\infty}\Phi_{m}(x)=0$.
+∎
+A.4 Proof of Corollary 4
+Proof.
+Set $N=\max\{n\mid a_{n}<1+\epsilon\}$.
+Define the sequence $\{b_{n}\}_{n=1}^{\infty}$ such that $b_{n}=a_{n}$ for $n\leq N$, and $b_{n}=1+\epsilon$ for $n>N$.
+The remainder of the proof is analogous to the proof of Proposition 3.
+∎
+A.5 Activation Distribution Based on Input Data Distribution
+Figure 8: The activation values in the 1000th layer, with $32$ nodes per hidden layer, were analyzed using the proposed weight initialization method with $\sigma_{z}$ values of $0.0003$, $0.015$, $0.3$, and $3$.
+The upper row shows results for $3000$ input samples drawn from a standard normal distribution, while the lower row presents results for samples drawn from a Beta distribution with parameters $a=2.0$ and $b=5.0$.
+Appendix B Classification Tasks
+In this section, we present experimental results for the benchmark classification datasets MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100 when employing the proposed weight initialization method.
+B.1 Width independence in Classification tasks
+[…]
+[…]
+[…]
+[…]
+[…]
+B.2 Non-uniform Hidden Layer Dimensions
+Tanh neural networks have been less widely employed compared to ReLU networks due to their higher computational complexity, susceptibility to the vanishing gradient problem, and the superior empirical performance of ReLU in various deep learning tasks.
+However, the recent success of PINNs using tanh neural networks has renewed interest in their application.
+This section compares the performance of four initialization methods on architectures that are generally challenging to train: (1) tanh activation with Xavier initialization, (2) tanh activation with the proposed initialization, (3) ReLU activation with He initialization + BN, and (4) ReLU activation with orthogonal initialization.
+The experiments are performed on FFNN and autoencoder models.
+FFNN.
+The experiment is performed using an FFNN with a structure in which hidden layers alternated between 16 and 4 nodes, repeated 50 times, and trained over 100 epochs.
+The results are shown in Figure 13 (a).
+Both Xavier and the proposed method successfully train the network, with the proposed method showing overall better performance.
+Due to its strong performance despite substantial variations in hidden layer sizes, the proposed method is further evaluated on autoencoders with large variations in layer sizes, as shown in Figure 13 (b).
+Autoencoder.
+The autoencoder architecture consists of an encoder and a decoder, both employing batch normalization and dropout (0.2) for regularization.
+The encoder compresses the input through layers of 512, 256, and 128 units before mapping to a 64-dimensional latent space.
+The decoder reconstructs the input by symmetrically expanding the latent space through layers of 128, 256, and 512 units, followed by an output layer with sigmoid activation.
+In Figure 13 (b), the model is trained on the MNIST dataset with a batch size of $256$, while in (c), it is trained on the FMNIST dataset with a batch size of $512$.
+(a) FFNN (MNIST)
+(b) Autoencoder (MNIST)
+(c) Autoencoder (FMNIST)
+Figure 13: (a) Validation loss for an FFNN with alternating hidden layers of $16$ and $4$ nodes, repeated $50$ times, comparing four methods: Tanh with Xavier initialization, Tanh with the proposed initialization, ReLU with He initialization + BN, and ReLU with orthogonal initialization.
+(b) Validation loss for an autoencoder with encoder-decoder layers of $512$, $256$, $128$, and $64$ units, comparing the same four methods.
+(c) Same as (b), but on the FMNIST dataset.
+Appendix C Physics-Informed Neural Networks
+In this section, we present additional experiments employing the proposed weight initialization method in Physics-Informed Neural Networks (PINNs).
+Specifically, we empirically analyze its relationship with different activation functions (Swish, ELU, Sigmoid, and ReLU) and its impact on various PDE problems (Burgers’ equation, Allen-Cahn equation, Poisson equation, and Diffusion equation).
+C.1 PDE Details
+Here we present the differential equations that we study in experiments.
+Allen-Cahn Equation.
+The diffusion coefficient is set to $d=0.01$.
+The initial condition is defined as $u(x,0)=x^{2}\cos(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=-1$ and $u(1,t)=-1$, applied over the time interval $t\in[0,1]$.
+The Allen-Cahn equation is expressed as
+$\frac{\partial u}{\partial t}-d\frac{\partial^{2}u}{\partial x^{2}}=-\frac{u^{3}+u}{d},$
+where $u(x,t)$ represents the solution, $d$ is the diffusion coefficient, and the nonlinear term $u^{3}-u$ models the phase separation dynamics.
+Burgers’ Equation.
+The Burgers’ equation, a viscosity coefficient of $\nu=0.01$ is employed.
+The initial condition is given by $u(x,0)=-\sin(\pi x)$ for $x\in[-1,1]$, with boundary conditions $u(-1,t)=0$ and $u(1,t)=0$ imposed for $t\in[0,1]$.
+The Burgers’ equation is expressed as
+$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}=\nu\frac{\partial^{2}u}{\partial x^{2}},$
+where $u(x,t)$ is the velocity field, and $\nu$ is the viscosity coefficient.
+Diffusion Equation.
+The diffusion equation includes a time-dependent source term and is defined over the spatial domain $x\in[-1,1]$ and temporal interval $t\in[0,1]$.
+The initial condition is specified as $u(x,0)=\sin(\pi x)$, with Dirichlet boundary conditions $u(-1,t)=0$ and $u(1,t)=0$.
+The diffusion equation is expressed as
+$\frac{\partial u}{\partial t}-\frac{\partial^{2}u}{\partial x^{2}}=e^{-t}\left(\sin(\pi x)-\pi^{2}\sin(\pi x)\right),$
+where $u(x,t)$ is the solution.
+Poisson Equation. The Poisson equation is defined over the spatial domain $x\in[0,1]$ and $y\in[0,1]$. The Poisson equation is expressed as
+$\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}=f(x,y),$
+where $u(x,y)$ is the solution, and $f(x,y)$ is the source term given by
+$f(x,y)=2\pi^{2}\sin(\pi x)\sin(\pi y).$
+C.2 Effect of activation function on PINNs
+We experimentally demonstrate that when employing the proposed weight initialization method, the absolute error between the exact solution and the PINN-predicted solution is smaller when using the tanh activation compared to ReLU, sigmoid, Swish (Ramachandran et al., 2017), and ELU (Clevert, 2015), as shown in Figure 14.
+The FFNN consists of 30 hidden layers (32 nodes each) and is trained for 300 iterations using Adam, followed by 300 iterations using L-BFGS.
+For the Burgers’ and diffusion equations, the PINN with tanh activation achieves the closest approximation to the exact solution.
+It is well established that PINNs with tanh activation effectively approximate solutions to the Burgers’ and diffusion equations.
+Given that the choice of activation function and weight initialization are dependent, this result provides valuable insights into the interaction between initialization methods and activation functions in PINNs..
+Figure 14: Absolute error for (upper row) the Burgers’ equation and (lower row) the diffusion equation with varying activation functions. Values outside the color bar range are depicted in white.
+Figure 15 presents experiments that employ the proposed initialization and Xavier initialization for PDE problems using the Swish activation function, which is widely used in PINNs.
+(a) Diffusion equation
+(b) Poisson equation
+(c) Burgers’ equation
+(d) Allen-Cahn equation
+Figure 15: PINN loss for a Swish FFNN with (a) $20$ hidden layers, each containing $32$ nodes, and (b), (c) $3$ hidden layers, each containing $3$ nodes, and (d) $10$ hidden layers, each containing $32$ nodes.
+C.3 $\sigma_{z}$ for Burgers’ Equation
 (a) Same dimension
 (b) Varying dimensions
-Figure 4: Here, ’STD’ refers to $\sigma_{z}$.
+Figure 16: Here, STD refers to $\sigma_{z}$.
 (a) shows the PINN loss for the Burgers’ equation, using an FFNN with 30 layers and 32 nodes in each hidden layer.
 (b) shows the PINN loss for an FFNN with 30 layers, where the hidden layers alternate between 64 and 32 nodes, repeated 15 times.
-Each experiment was repeated 10 times with different random seeds.
-B.2 Width independence in Classification tasks
-Figure 5: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers.
-All models were trained for 20 epochs on the MNIST dataset, with 10 different random seeds.
-Figure 6: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers.
-All models were trained for 20 epochs on the Fashion MNIST dataset, with 10 different random seeds.
-Figure 7: Validation accuracy and loss are presented for tanh FFNNs with varying numbers of nodes $(2,8,32,128)$, each with 20 hidden layers.
-All models were trained for 20 epochs on the CIFAR-10 dataset, with 10 different random seeds.
+Here, we present the PINN loss for the Burgers’ equation as a function of the initialization hyperparameter $\sigma_{z}$.
+As shown in Figure 16, the experimental results demonstrate that $\sigma_{z}$ significantly impacts training.
+Therefore, selecting an appropriate $\sigma_{z}$ is crucial for optimizing performance.
+C.4 Absolute error for Burgers’ equation
+Figure 17 demonstrates that increasing the number of collocation points allows PINNs with the proposed initialization to predict solutions with lower absolute error compared to those using Xavier initialization.
+The proposed method is experimentally demonstrated to achieve faster convergence speed and higher data efficiency.
+Figure 17: Absolute error between the exact solution and the PINN-predicted solution for the Burgers’ equation with varying numbers of collocation points (3000, 10000, 20000, 50000) using (upper row) Xavier and (lower row) the proposed initialization.
+The FFNN has 30 hidden layers ($32$ nodes each) and is trained for $300$ iterations using Adam followed by $300$ iterations using L-BFGS.
+Appendix D Examples of the matrix $\mathbf{D}^{\ell}$
+An example of $\mathbf{D}^{\ell}$ from the initialization methodology proposed in Section 3.2 is presented.
+In the heatmap of Figure 18, white represents the value 00, while black represents the value $1$.
+Figure 18: Examples of the matrix $\mathbf{D}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ in Section 3.2 with $N_{\ell}=4,N_{\ell-1}=8$ (left), $N_{\ell}=4,N_{\ell-1}=4$ (middle), and $N_{\ell}=8,N_{\ell-1}=4$ (right), respectively.